### v1

# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

Abstract
Federated Domain-specific Instruction Tuning (FedDIT) leverages a few cross-client private data and server-side public data for instruction augmentation, enhancing model performance in specific domains. While the factors affecting FedDIT remain unclear and existing instruction augmentation methods mainly focus on the centralized setting without considering the distributed environment. Firstly, our experiments show that cross-client domain coverage, rather than data heterogeneity, drives model performance in FedDIT. Thus, we propose FedDCA, which maximizes domain coverage through greedy client center selection and retrieval-based augmentation. To reduce client-side computation, FedDCA uses heterogeneous encoders with server-side feature alignment. Extensive experiments across four domains (code, medical, financial, and mathematical) validate the effectiveness of both methods. Additionally, we explore the privacy protection against memory extraction attacks with various amounts of public data and results show that there is no significant correlation between the amount of public data and the privacy-preserving capability. However, as the fine-tuning round increases, the risk of privacy leakage reduces or converges.
1 Introduction
Table 1: Performance(%) of different augmentation settings on each domain, which are conducted via FedAvg protocol with 10 clients. Specifically, the zero-shot directly inferences without fine-tuning, while base data utilizes only the client’s local data for FedDIT. Also, we compare FedDCA with other two augmentation strategies: random sampling and direct retrieval (described in Appendix A.3), respectively.
Metric Zero-shot Base Data Random Sampling Direct Retrieval FedDCA
HumanEval/Pass@1 29.88 39.03 32.93 34.14 36.58
MMLU-Med/Acc 70.60 68.40 71.30 72.20 74.50
FPB/Acc 55.94 58.25 64.19 66.31 67.24
FiQA/Acc 18.54 14.18 13.09 19.11 35.27
TFNS/Acc 59.21 66.62 65.53 67.62 73.32
GSM8K/Exact Match 23.27 47.46 47.38 50.87 52.46
Avg. 42.91 48.99 49.07 51.71 56.56
Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to enable collaborative training of large language models (LLM) in a distributed environment without the need to share private data (McMahan et al. (2017); Ye et al. (2024b)). This approach involves exchanging model parameters among distributed data holders, striking a delicate balance between preserving privacy and optimizing model efficiency. Although many FedIT frameworks have been established(Ye et al. (2024b); Kuang et al. (2023); Zhang et al. (2024a)), they have not deeply explored the practical issues that FedDIT might encounter in real-world scenarios. For example, FedIT usually requires sufficient instruction data for fine-tuning, which is difficult to meet in domain-specific fine-tuning scenarios (Zhang et al. (2024d)).
In this study, we explore Federated Domain-specific Instruction Tuning (FedDIT), a novel approach within the FL paradigm that focuses on enhancing the performance of LLMs in specific domains. Unlike the general FedIT, which seeks to improve model performance across a variety of tasks without considering local data shortage, FedDIT faces the challenge of clients having only a few local domain-specific data. To overcome this, FedDIT leverages a server-hosted public dataset that spans multiple domains to enrich the local data through the specific instruction augmentation strategy. This strategic enrichment is essential for achieving effective instruction tuning and needs to be carefully designed to prevent performance degradation. Except for the code domain, which is more about following a certain paradigm. As shown in Table 1, when clients attempt to do FedDIT only based on their local data, although some high-quality in-domain local data may be available, the limited scale restricts effective instruction tuning. Consequently, this sometimes leads to performance degradation, as indicated by the underlined values. Overall, the goal of FedDIT is to develop a domain-specific LLM that utilizes collaborative training and instruction augmentation with public data while preserving client privacy, ensuring that the model is adept at handling tasks within its targeted domain.
Additionally, the factors affecting FedDIT are still unclear. What’s worse, augmenting instructions may make the results even more unpredictable and confused about how to improve. Shepherd(Zhang et al. (2024a)) approaches this problem from the perspective of heterogeneity, constructing heterogeneity based on the topic of general instruction datasets. The study demonstrates that, for general FedIT, unlike the consensus of traditional FL, heterogeneity has a positive effect. By aggregating diverse instructions from clients, the diversity increases, thereby enhancing the model’s adaptability to various tasks. However, it just scratches the surface and does not explore issues in FedDIT.
Going one step further, we conduct experiments to unveil a new key finding (Appendix A.2): there is no linear correlation between the degree of non-independent and identically distributed (non-iid) and the performance of LLM for FedDIT. Inspired by Explore-Instruct(Wan et al. (2023)), we first define the domain coverage metric and investigate the effect of domain coverage. Results show that the cross-client domain coverage impacts the model’s performance significantly.
To maximize the cross-client domain coverage without compromising client data privacy, based on INSTA(Lee et al. (2024)), we propose a novel FedDIT algorithm, called FedDCA (Federated Instruction Tuning of LLMs with Domain Coverage Augmentation), which performs greedy client center selection and instruction augmentation via dense retrieval on the server side. The core idea of FedDCA lies in greedy client center selection, which chooses client centers that enhance the diversity and coverage of the domain-specific characteristics within the public dataset. By concentrating on maximizing domain coverage at each step, the algorithm efficiently builds an augmented train set that boosts the model’s learning and generalization capabilities, resulting in superior performance on domain-specific tasks. Furthermore, to reduce the computation overhead on the client side, we propose FedDCA with heterogeneous encoders. For feature alignment, we train a projector on the server side using public data through contrastive learning.
We demonstrate the effectiveness of FedDCA by conducting extensive experiments across four domains: code, medical, financial, and mathematical, and comparing against multiple baselines, which can be categorized into two types: unaugmented and augmented methods. For the unaugmented setting, we compare our method with FedIT, which composes of four orthodox FL methods, including FedAvg(McMahan et al. (2017)), FedProx(Li et al. (2020)), SCAFFOLD(pmlr-v119-karimireddy20a), and FedAvgM(Hsu et al. (2019)). For the augmented setting, we compare our method with random sampling, direct retrieval, LESS(Xia et al. (2024)), and Self-Instruction(Wang et al. (2023c)). Additionally, we report FedDCA’s performance with different FL strategies. Furthermore, we compare the computational efficiency of FedDCA and FedDCA on the client side. To illustrate the impact of different retrieval volumes on client data privacy, we conduct experiments against memory extraction attacks. Results indicate that as the number of training rounds increases, the method only based on the client’s local data memorizes more privacy information while the privacy leakage risk gradually decreases or converges for the augmented setting.
The main contribution is as follows:
  • We unveil a key finding: For FedDIT, there is no linear correlation between the degree of heterogeneity and the model’s performance. Instead, cross-client domain coverage impacts the model’s performance significantly, which is presented in Appendix A.2.
  • We propose a novel FedDIT algorithm (Section 4) called FedDCA to maximize the cross-client domain coverage through greedy client center selection and then perform retrieval-based instruction augmentation on the server. To further reduce the computational overhead on the client side, we propose FedDCA with a heterogeneous encoder setting. For feature alignment, we train a projector on the server side.
  • With extensive experiments (Section 5), we show the effectiveness of FedDCA and FedDCA. Furthermore, we explore privacy preservation against memory extraction attacks and conduct experiments based on various public data amounts. Results indicate that the privacy-preserving capability is not significantly correlated with the amount of public data, but the risk of privacy leakage decreases or converges as the amount of fine-tuning rounds increases.
2 Preliminaries
Federated domain-specific instruction tuning (FedDIT) aims to leverage cross-client private domain-specific instruction data and utilize the multi-domain public data on the server to achieve instruction augmentation, collaboratively enhancing the model’s performance in specific domains. Suppose $N$ distributed clients with their local private data $D_{k}^{l}$. The server has a public dataset $D^{p}$ consisting of multiple domains and is responsible for data augmentation based on different strategies and aggregation model parameters from clients. The training process follows standard FL protocol(McMahan et al. (2017)). For computational efficiency, we adopt LoRA(Hu et al. (2021)) as the fine-tuning method, which involves tuning additional parameters $\Delta\phi$ while keeping the pre-trained LLM’s parameters $\phi$ frozen. In the initial training stage, the server dispatches the pre-trained LLM parameters $\phi$ and the initialized LoRA parameters $\Delta\phi$ to each client for training $\mathcal{R}$ rounds. In the $t$-th round, the server sends the aggregated LoRA parameters $\Delta\phi^{t}$ to clients. Clients use $\Delta\phi^{t}$ to update their local LoRA parameters $\Delta\phi_{k}^{t}$ and perform instruction tuning based on augmented instruction datasets, then send the updated LoRA parameters $\Delta\phi_{k}^{t+1}$ back to the server. The server receives the LoRA parameters $\{\Delta\phi_{k}^{t+1}\mid k=1,\dots,N\}$ from each client and aggregates them to obtain the global LoRA parameters $\Delta\phi^{t+1}$ for the next round.
3 Problem Formulation
For better understanding, we list the frequently used notation in Appendix A.11. The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data(Ye et al. (2024b); Zhang et al. (2024d)). Under the FL framework, suppose we have $N$ clients, where each client $c_{k}$ has a local dataset $D_{k}^{l}$ with its size $N_{k}^{l}$, and an augmented dataset $D_{k}^{g}$ from the server public dataset $D^{p}$, respectively. Due to constraints such as memory, computational overhead, and maximum tolerated training time, client $c_{k}$ can accept at most $N_{k}^{p}$ public instructions. Denote the augmentation strategy as $\Lambda$, through which the server performs instruction augmentation on the public dataset. If $\Lambda$ is null, it denotes clients only perform FedDIT based only on their local private data, which may cause performance degradation. Otherwise, $\Lambda$ could be categorized into two types: 1) Focusing only on its own local data distribution. 2) Taking into account the cross-client data distribution. In conclusion, the global objective of FedDIT is defined as follows:
$\operatorname*{arg\,min}_{\Delta\phi}{\left\{F(\phi,\Delta\phi)\triangleq\sum_{k=1}^{N}p_{k}\left((1-\alpha_{k})F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{l}\right)+\alpha_{k}F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{g}\right)\right)\right\}},$ (1)
where the first term and second term denote the accumulated fine-tuning loss computed on the client $c_{k}$’s local instructions $D_{k}^{l}$ and the augmented public instructions $D_{k}^{g}$ from the server respectively. $p_{k}$ is the weight of the $k$-th client, which is determined by the ratio of $k$-th client’s data size to all clients’ total data size. $\alpha_{k}$ is the ratio of the public data amount of its augmented instruction dataset $D_{k}$, which is computed as $\frac{N_{k}^{p}}{N_{k}^{l}+N_{k}^{p}}$. Eq.1 minimizes the summed empirical loss across clients’ augmented instructions to pursue the in-domain utility of the obtained global model. Denote $P$ as the model parameters and $\mathcal{D}$ as a specific dataset, then the client $c_{k}$’s empirical loss $F_{k}(\phi,\Delta\phi_{k};\mathcal{D})$ base on $\mathcal{D}$ is calculated as:
$F_{k}(\phi,\Delta\phi_{k};\mathcal{D})\triangleq\frac{1}{|\mathcal{D}|}\sum_{j=1}^{|\mathcal{D}|}l(P_{\phi+\Delta\phi_{k}};x_{j}),$ (2)
where $x_{j}\in D,\forall j\in\{1,2,\dots,|D|\}$. The instruction tuning loss $l(\cdot;\cdot)$ on a sample $(x,y)$ is defined as $-\sum_{t=1}^{\left|y\right|}\log\left(w\left(y_{t}|x,y_{<t}\right)\right)$, where $x$ is the formated instruction with Alpaca instruction template [^1] https://crfm.stanford.edu/2023/03/13/alpaca.html and $y$ is the corressponding response.
In contrast to the problem formulations in many prior works on FedIT, the primary distinction in the formulation of FedDIT lies in acknowledging the lack of in-domain instructions across clients. FedDIT establishes a public multi-domain dataset on the server for domain-specific instruction augmentation by a specific sampling strategy $\Lambda$. Therefore, each client only needs to hold a few high-quality domain-specific private data to collaborate with other clients to obtain a strong domain-specific LLM.
4 Method
We propose Federated Instruction Tuning of LLMs with Domain Coverage Augmentation (FedDCA), which enhances domain coverage through greedy client center selection and performs retrieval-based instruction augmentation to obtain a LLM that performs well on domain-specific tasks. We first give the overview FedDCA (Section 4.1) and define the domain coverage metric (Section 4.2), then show details of the greedy client center selection algorithm (Section 4.3) and heterogeneous encoder method FedDCA (Section 4.4).
4.1 Overview
Figure 1: Overview of FedDCA, which consists of three stages: 1) The client $c_{k}$ performs local instructions clustering and sends cluster centers $\mathcal{C}_{k}$ to the server. 2) The server first does greedy client center selection to maximize domain coverage and performs client-center-based domain data retrieval, then sends the augmented instructions $D_{k}^{g}$ to the client $c_{k}$. 3) Clients fine-tune the LLM collaboratively and the LoRA parameters $\Delta\phi$ are exchanged between clients and the server.
As shown in Figure 1, FedDCA consists of three stages, in the first stage the client $c_{k}$ clusters its local instructions and sends the cluster centers $\mathcal{C}_{k}$ to the server. Then the server performs greedy client center selection to obtain client centers that maximize the domain coverage, then performs client-center-based retrieval for domain-specific instruction augmentation. Additionally, the third stage follows the standard FedIT procedure(Ye et al. (2024b)), in which the server sends the frozen pre-trained LLM parameters $\phi$ and the initialized LoRA parameters $\Delta\phi$ to clients and then $\Delta\phi_{k}$ is exchanged between the client and the server.
The novel design of FedDCA lies in the second stage before training. As shown in Algorithm 1, the core idea is to use a greedy client center selection algorithm to obtain client centers that maximize cross-client domain coverage, and then use these client centers to retrieve instructions on the public dataset. To further reduce the computational overhead of encoding on clients, we propose a heterogeneous encoder method called FedDCA. Specifically, the client uses a small encoder locally, while the server uses a larger encoder for accurate semantics. To align features with different dimensions, we introduce a projector $w_{p}$ on the server side, which is obtained through contrastive learning.
Algorithm 1 FedDCA: greedy client center selection & domain data retrieval
1:Parameters: Number of clusters $\xi$; Encoder model $w_{enc}$; Client local datasets $D=\{D_{1},D_{2},\ldots,D_{N}\}$; Public dataset $D^{p}$; Similarity score threshold $\alpha$; Number of public data samples retrieved per client $N_{k}^{p}$; Client centers $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{N}\}$; Pre-encoded public instruction embeddings $\mathcal{E}$.
2:for $i\in\{1,2,\ldots,N\}$ do
3:     $\mathcal{E}^{\prime}\leftarrow\{w_{enc}(x)\mid x\in D_{i,instruction}^{l}\}$
4:     $\mathcal{C}_{i},S_{i}\leftarrow\text{k-means}(\xi).\text{fit}(\mathcal{E^{\prime}})$triangleright Cluster local instructions, return cluster centers $\mathcal{C}_{i}$ and sizes $S_{i}$
5:end for
6:Send $\{\mathcal{C}_{i},S_{i}\mid i\in\{1,2,\ldots,N\}\}$ to the server for the greedy client center selection
7:$\mathcal{P}\leftarrow\{\mathcal{C}_{0,k}\ \mid k=\arg\max(S_{0})\}$triangleright Initialize the client center set
8:for $i\in\{1,2,\ldots,N-1\}$ do $\triangleright$ Greedy client center selection
9:     $\mathcal{S}\leftarrow\sum(\mathcal{C}_{i}\cdot\mathcal{P}^{T},\text{dim}=-1)$triangleright Compute summed similarity score of each cluster center
10:     $\mathcal{I}\leftarrow(N-i)\text{-}\arg\text{sort}(\mathcal{S})$triangleright Filter $i$ cluster centers close to the client center set $\mathcal{P}$
11:     $\mathcal{I^{\prime}}\leftarrow\arg\text{sort}(-S_{i})$triangleright Sort cluster center by cluster size in descending order
12:     $j\leftarrow\text{first element of }\mathcal{I}\cap\mathcal{I}^{\prime}$triangleright Selected cluster center $\mathcal{C}_{i,j}$
13:     $\mathcal{P}\leftarrow\mathcal{P}\cup\mathcal{C}_{i,j}$triangleright Update the client center set
14:end for
15:for $i\in\{1,2,\ldots,N\}$ do $\triangleright$ Client center based domain data retrieval
16:     $\mathcal{S}\leftarrow\mathcal{P}_{i}\cdot\mathcal{E}^{T}$triangleright Compute similarity score between $\mathcal{P}_{i}$ and $\mathcal{E}$
17:     $\mathcal{S^{\prime}}\leftarrow\{s\mid s\in\mathcal{S},s<\alpha\}$triangleright Filter instructions with similarity score larger than $\alpha$
18:     indices of $\mathcal{I}\leftarrow\text{indices of }N_{k}^{p}\text{-top in }\mathcal{S^{\prime}}$
19:     $D_{i}^{g}\leftarrow\{D^{p}_{j}\mid j\in\mathcal{I}\}$
20:     $D_{i}\leftarrow D_{i}^{l}\cup D_{i}^{g}$triangleright Obtain the augmented instruction dataset $D_{i}$
21:end for
4.2 Domain Coverage
As domain coverage directly affects the in-domain performance of the LLM (Appendix A.2), FedDCA aims to maximize the domain coverage of the cross-client data distribution $\cup_{i=1}^{N}D_{i}$ respect to the in-domain data distribution $D^{d}$, which is defined in Eq. 5. However, as public data is unreachable for clients and clients can not send the local data to the server for privacy, directly finding a cross-client dataset that maximizes the domain coverage in Eq. 5 is unrealistic. To find a proper client center set $\mathcal{P}$ that attempts to maximize the domain coverage $d(D^{d},P)$ and use $\mathcal{P}$ to perform instruction retrieval on the public data is an approximation problem.
Suppose there exists a metric space $\mathcal{X}$, a set of cross-client local instruction embeddings $\mathcal{E}\in\mathcal{X}$, and a set of cluster centers $\mathcal{P}\in\mathcal{X}$. Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm(Vardakas & Likas (2023)). As in the federated setting, communication cost is always considered. Thus, we need to find a $\mathcal{P}$ that maximizes the domain coverage while considering the communication cost, we formulate the optimization problem as follows:
$\operatorname*{arg\,min}_{\mathcal{P}}\left\{\sum_{i=1}^{N}|\mathcal{P}_{i}|+\sum_{d\in D^{d}}\left(\text{min}_{p\in\mathcal{P}}sim(d,p)\right)\right\},$ (3)
[…]
4.3 Greedy Client Center Selection & Domain Data Retrieval
As described in Section 4.2, the goal of FedDCA is to find a proper client center set $\mathcal{P}$ that maximizes the domain coverage $d(D^{d},\mathcal{P})$, as defined in Eq. 3. For computational efficiency, we propose a greedy algorithm as shown in Algorithm 1 and Figure 2 to solve this problem in polynomial time and obtain a sub-optimal solution. Given the cluster centers $\{\mathcal{C}_{i},i=1,2,\dots,N\}$ received from each client, which are obtained by clustering local instructions. FedDCA on the server consists of two main steps: greedy client center selection and client-center-based domain data retrieval.
Greedy client center selection. We will consider this problem from two aspects: 1) Select a client center that can represent the distribution of the local data. 2) To maximize the cross-client domain coverage, we filter client centers that are close to the previously selected client centers. Firstly, we randomly choose a client and select the largest cluster center as its center and be the initial of the client center set $\mathcal{P}$. Then we iteratively select the client center $\mathcal{P}_{k}$ of the top cluster size, while avoiding selecting the cluster center close to the previously selected client center to maximize the domain coverage of $\mathcal{P}$. Specifically, for the $i$-th iteration, we filter $i$-top cluster centers that have the largest summed similarity with previously selected client centers and select the center of the rest largest cluster as the $i$-th client center. This procedure is repeated until have selected $N$ client centers.
Domain data retrieval. For each client center $\mathcal{P}_{k}$, the server performs dense retrieval (detailed in Appendix A.3) on public dataset $D^{p}$ to get the top-$N_{k}^{p}$ similar public instructions, then sends $\{D_{1}^{g},\dots,D_{N}^{g}\}$ to each clients. Specifically, to avoid the overlap between public data and local private data, we set a threshold $\alpha$ to filter the public instructions that have a similarity score larger than $\alpha$ with the client center.
In summary, FedDCA can be seen as a method for establishing a comprehensive training set that captures the essence of the domain and covers its distribution. Through the iterative selection of client centers that provide unique coverage, the algorithm ensures that the selected centers are representative of the domain and achieve great domain coverage. In addition, domain data retrieval exposes the model to a broader range of in-domain features, facilitating a more robust understanding of domain-specific tasks. This not only enriches the training data but also reduces the risks of overfitting to the limited local data available to each client, which ends up with performance degradation. By leveraging diverse examples across clients, the model can better generalize its learning, leading to enhanced performance on various tasks within the domain.
Figure 2: Greedy client center selection iteratively selects the client center in each step, which synthetically considers both the representativeness of the cluster center and its distance from the previously selected client center set $\mathcal{P}$ to maximize the cross-client domain coverage.
4.4 Heterogeneous Encoder with Feature Alignment
Due to the limited computational resources on the client side, we propose a heterogeneous encoder method FedDCA to reduce the client’s computational overhead by using a small encoder $\omega$ on the client side and a larger encoder $\omega^{\prime}$ on the server side. However, the output dimension of heterogeneous encoders may not be consistent. Therefore, we introduce a projector $w_{p}$ on the server for feature alignment, as shown in Appendix A.4, we use contrastive learning(Khosla et al. (2020)) and train $w_{p}$ on the public instructions. Specifically, $w_{p}$ is composed of two fully connected layers and a ReLU activation layer in between, projecting $\omega$’s dimension to the output dimension of $\omega^{\prime}$. For $\forall{x_{i}}\in D^{p}$, where $i\in\{1,2,\dots,\mathcal{N}\}$, $\omega$ outputs embedding $h_{i}$, $\omega^{\prime}$ outputs embedding $h_{i}^{\prime}$ and $w_{p}$ outputs embedding $\varphi_{i}$. For input $x_{i}$, the positive sample pair is $(h_{i}^{\prime},\varphi_{i})$ and negative sample pairs are $\{\varphi_{j},j\neq i\}$. Let the batch size be $B$, then the training objective is defined as follows:
$\mathcal{L}=-\sum_{i=1}^{B}\log\frac{\exp\left(\text{sim}(\varphi_{i},h^{\prime}_{i})/\tau\right)}{\sum_{j=1}^{B}\mathbb{I}_{[j\neq i]}\exp\left(\text{sim}(\varphi_{i},h^{\prime}_{j})/\tau\right)},$ (4)
where $\text{sim}(\cdot,\cdot)$ denotes the cosine similarity and $\tau$ denotes the temperature parameter.
Using a projector $w_{p}$ to achieve feature alignment is an effective strategy in FL for scenarios where clients and the server utilize encoders of different sizes or capacities. The projector maps the lower-dimensional feature representations $\mathcal{H}$ from the client-side encoder to the higher-dimensional space $\mathcal{H}^{\prime}$ of the server-side encoder. By presenting the projector with positive and negative pairs, $w_{p}$ learns to adjust the projections such that the feature representations of similar items are close to each other, while those of dissimilar items are distant. In conclusion, the heterogenous encoder setting allows the system to be more scalable and flexible based on different client capacities. As a result, the projector helps in creating a unified feature space where representations from diverse sources can be combined, and utilized collaboratively.
4.5 Discussions
Computation. Through greedy client center selection, FedDCA solves the optimization problem defined in Eq. 3 in the polynomial time and quite efficient. On the client side, the k-means algorithm(Vardakas & Likas (2023)) is $\mathcal{O}(|D_{k}^{l}|)$. On the server side, as mentioned in Section 4.2, the time complexity of greedy client center selection is $\mathcal{O}\left(N\left(N\xi+\xi\log\xi\right)\right)$. Specifically, for each client, the similarity computation between selected client center set $\mathcal{P}$ and client $c_{k}$’s cluster $\mathcal{C}_{k}$ is $\mathcal{O}\left(N\xi\right)$, and the sorting process is $\mathcal{O}\left(\xi\log\xi\right)$. For domain data retrieval, given that the sorting algorithm is $\mathcal{O}\left(\mathcal{N}\log\mathcal{N}\right)$, where $\mathcal{N}$ is the size of the public dataset, thus the time complexity is $\mathcal{O}\left(N(\mathcal{N}\log\mathcal{N})\right)$. In addition, the hetergenous encoder setting FedDCA further reduces the computation overhead on the client side.
Communication. In the domain instruction augmentation stage, each client sends $\xi$ cluster centers to the server. Next, the server performs greedy client center selection and domain data retrieval, then sends retrieved public data $D_{k}^{g}$ to the client $c_{k}$, which size is $N_{k}^{p}$. In addition, in the fine-tuning stage, FedDCA follows the standard FedIT procedure, which exchanges the LoRA parameters $\Delta\phi_{k}$ between clients and the server. Specifically, for Llama3-8B model, the number of trainable LoRA parameters is just 13.6 million. Compared with the number of the frozen pretrained LLM parameters, the communication for LoRA tuning is quite efficient.
Privacy. The proposed method FedDCA does not violate the client’s privacy. Compared with other FedIT methods(Zhang et al. (2024d); Ye et al. (2024b)), the difference lies in the greedy client center selection. In this stage, the client only uploads the cluster center to the server, which is the average of embeddings to its cluster. In addition, the potential privacy leakage can be further avoided through homomorphic encryption(Acar et al. (2018)), which allows the server directly computates on ciphertext for matrix multiplication for dense retrieval.
5 Experiments
To show the effectiveness of FedDCA, we conduct extensive experiments on various domains and compare FedDCA with several baselines. This section is organized as follows: First we give details of experimental setup in Section 5.1, then we report the performance of FedDCA and compare it with several baselines in Section 5.2, and next we evaluate the computational cost between the FedDCA and FedDCA in Section 5.3. Finally, we do privacy evaluations of FedDCA in Section 5.4. For more details and results please see in Appendix A.
5.1 Experimental Setup
Dataset and Evaluation Metrics. To evaluate the performance of FedDCA, we conduct experiments on the four domains: code, medical, financial and mathematcial. The detail of constructing the public dataset is described in Appendix A.5. We choose HumanEval (code, H-Eval for short)(Chen et al. (2021)), MMLU-Med (medical, abbreviated as M-Med)(Hendrycks et al. (2021)), FPB (financial), FiQA (financial), TFNS (financial)(Yang et al. (2023a)), GSM8K (mathematical)(Cobbe et al. (2021)) for evaluation. HumanEval is evaluated using Pass@1. For MMLU-Med, FPB, FiQA and TFNS, we use accuracy as evaluation metric and exact match (EM) is used for GSM8K. Please see Appendix A.5 for more dataset details.
Baselines. We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented. Zero-shot inference and FedIT, which composes of four widely used FL methods, including FedAvg(McMahan et al. (2017)), FedProx(Li et al. (2020)), SCAFFOLD(pmlr-v119-karimireddy20a) and FedAvgM(Hsu et al. (2019)). 2) Augmented. Random sampling, direct retrieval, LESS(Xia et al. (2024)), Self-Instruct(Wang et al. (2023c)). Additionally, we report FedDCA’s performance with different FL strategies in Table 2. Specifically, zero-shot inference shows the performance of the pre-trained LLM without FedDIT, which gives the lower performance bound. Direct retrieval is described in Appendix A.3. Self-Instruct augments instruction from a generative way through GPT-3.5-turbo(Sun et al. (2023)). Base on the prompt provided by Self-Instruct, we define the prompts for generating instructions and responses as shown in Appendix A.7.
5.2 Performance Analysis
Performance. We evaluate the performance of FedDCA and compare it with several baselines on four domains (see training details in Appendix A.6). As shown in Table 2, FedDCA outperforms the other nine baselines in all domains. In particular, FedDCA+FedProx achieves the best performance in most domains, with a substantial improvement from at least 5.61% to the maximum of 14.63% over other baselines on the average performance. In addition, to reduce the computation overhead of clients, FedDCA attempts to utilize heterogeneous encoders. We see that although FedDCA has a performance drop of 1.67% than FedDCA, it still outperforms other baselines in the average performance. Furthermore, we report FedDCA’s performance with different FL strategies, we can see that no FL methods can keep the leading position in all domains. In specific, FedProx and SCAFFOLD FL strategies perform better in the average performance. Overall, the result shows the effectiveness of FedDCA and FedDCA.
Table 2: Performance(%) of FedDCA, FedDCA and other nine baselines on various domains. We report FedDCA with different FL strategies in the last four rows and FedDCA+FedProx achieves the best average performance across four domains.
Method H-Eval M-Med FPB FiQA TFNS GSM8K Avg.
Zero-shot 29.88 70.60 55.94 18.54 59.21 23.27 42.91
FedAvg 39.03 68.40 58.25 14.18 66.62 47.46 48.99
FedProx 37.20 69.10 56.51 14.90 66.45 47.15 48.55
SCAFFOLD 37.80 70.20 62.71 15.27 66.49 49.27 50.29
FedAvgM 32.32 64.70 68.14 29.27 70.32 46.85 51.93
Random Sampling 32.93 71.30 64.19 13.09 65.53 47.38 49.07
Direct Retrieval 34.14 72.20 66.31 19.11 67.62 50.87 51.71
LESS 28.04 71.00 60.56 16.00 61.14 43.13 46.65
Self-Instruct 32.92 71.90 59.73 20.67 66.54 50.79 50.43
FedDCA 34.75 73.30 67.10 30.54 71.01 51.55 54.71
FedDCA+FedAvg 36.58 74.50 67.24 35.27 73.32 52.46 56.56
FedDCA+FedProx 32.92 72.40 72.93 38.18 77.55 51.25 57.54
FedDCA+SCAFFOLD 39.87 73.20 72.68 33.09 75.50 50.26 57.29
FedDCA+FedAvgM 33.53 68.90 71.45 31.45 72.52 49.76 54.60
Furthermore, Table 2 can be divided into two parts: unaugmented methods and augmented methods. We can observe that FedIT methods perform well in the code domain, even better than most augmented methods. This could be attributed to two reasons: 1) The code domain is more about following a certain paradigm. 2) As the local data is few, so compared with the same epoch and batch size, the unaugmented methods learn the same data more times, which is kind of not a fair setting. However, FedDCA+SCAFFOLD still surpasses the best baseline in the code domain, which further demonstrates the effectiveness of FedDCA.
Efficiency. We see that direct retrieval is the best baseline in the average performance. However, it does not consider the cross-client domain coverage, resulting in an overlapping retrieval data distribution. Self-Instruct(Wang et al. (2023c)) represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process. On the other hand, FewFedPIT(Zhang et al. (2024d)) attempts to augment local data by leveraging the pre-trained Llama2 model(Li & Flanigan (2023)) to perform self-instruct manner instruction generation during training. However, there exist two concerns: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation. 2) Generating instructions locally is very costly in terms of both time and computational resources. Overall, generating instructions through a self-instruct manner is not a satisfactory method for the FedDIT scenario. LESS(Xia et al. (2024)) represents the augmentation methods through gradient feature retrieval, however, the performance of LESS is not satisfactory. What’s worse, it assumes success to the validation set which is not always available and requires a warmup process on the public dataset first and then computes the gradient of both the public data and the validation set, which is computationally expensive when the size of the public dataset is enormous. In conclusion, FedDCA achieves better performance while requiring lower computational resources and a more relaxed training condition, which shows the efficiency of FedDCA.
Table 3: Domain coverage of FedDCA, FedDCA and other baselines on four domains, where FedIT methods only use local data without domain augmentation.
Method Code Med. Fin. Math. Avg.
FedIT 0.8126 0.6990 0.8529 0.7871 0.7879
Random Sampling 0.8512 0.7940 0.9196 0.8651 0.8575
Direct Retrieval 0.9396 0.8830 0.9293 0.8967 0.9122
LESS 0.8509 0.7737 0.8917 0.8352 0.8379
Self-Instruct 0.8966 0.8586 0.9015 0.8811 0.8845
FedDCA 0.9532 0.8972 0.9538 0.9096 0.9285
FedDCA 0.9766 0.9348 0.9815 0.9320 0.9562
Domain coverage. The domain coverage of each method is shown in Table 3. Also, we show the visualization of the domain coverage across different instruction augmentation methods in Appendix A.10. For the augmented methods, it can be observed that the higher domain coverage in a particular domain correlates with the better performance of the method in that domain. Specifically, given that FedDCA aims to maximize the domain coverage through greedy client center selection, therefore it achieves the highest domain coverage in all four domains, which surpasses other baselines from 4.82% to 21.36% average relative improvement. In addition, as FedDCA utilizes a smaller encoder on the client side for computational efficiency, it sacrifices some semantic precision, resulting in a 2.89% averaged domain coverage drop in relative. However, FedDCA still surpasses the best baseline by 1.79% on average. In brief, the result proves the effectiveness of FedDCA and FedDCA in domain coverage augmentation.
5.3 Computation Analysis
Table 4: Computational cost comparison between FedDCA and FedDCA. We report the encoding time on the public dataset and the model size of each encoder.
Method Encoding Time Model Size
FedDCA 15 min 46 s 335 M
FedDCA 5 min 02 s 22.7 M
Heterogenous encoder setting FedDCA allows the system to be more scalable and flexible based on different client capacities. To show the computational efficiency of FedDCA compared with FedDCA, we compare these two methods from the following aspects: 1) Encoder model size. 2) Encoding time overhead. We first give the evaluation setup of computation analysis.
Evaluation setup. The encoders used for FedDCA and FedDCA on the client side are bge-large-en-v1.5 and all-MiniLM-L6-v2 respectively (detailed in Appendix A.6). Then we show the model size and the time overhead of encoding the public instructions in Table 4
We can observe that in comparison with FedDCA, there is a large computational and time overhead advantage with FedDCA, while achieves an acceptable performance (shown in Table 2), which is accomplished by implementing a learnable projector on the server side. This projector maps the lower-dimensional client-side feature to the higher-dimensional space of the server-side encoder output. Consequently, the projector facilitates the creation of a unified feature space where representations from diverse sources can be integrated and used collaboratively.
5.4 Privacy Analysis
Next, we evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks(Carlini et al. (2021); Zhang et al. (2024c)), which utilizes the autoregression nature of LLM to extract information from the memory(Xu et al. (2024)).
Table 5: Privacy-preserving capability against memory extraction attack with different amounts of augmented public data. We report the average ROUGE-L score of each setting in four domains.
Retrieval Num Code Med. Fin. Math.
Base Data 0.2338 0.1579 0.1527 0.1718
1000 0.2228 0.1505 0.1451 0.1635
2000 0.2226 0.1531 0.1445 0.1627
3000 0.2195 0.1513 0.1454 0.1555
4000 0.2162 0.1434 0.1438 0.1631
5000 0.2119 0.1501 0.1445 0.1647
Evaluation setup. We focus on a single client’s fine-tuning process, which has 100 local data and adopts FedDCA to do the instruction augmentation with different amounts of public data from 1000 to 5000. We set up 10 clients with full participation for 10 rounds. Specifically, we record every round’s ROUGE-L score(Lin (2004)) of the client $c_{0}$. We use the designed prompt to extract the instruction that LLM memorizes, which is shown in Appdix A.8. Specifically, for each setting, we repeat the memory extraction 100 times and report the average ROUGE-L score computed based on each generated instruction and all the local data’s instructions to evaluate the privacy-preserving capability of different amounts of public data.
Figure 3: ROUGE-L score of each pair of base-data-only and augmented fine-tuning in four domains per round, where different monochromatic represents each domain.
Results. We report the average ROUGE-L score of base-data-only fine-tuning and various amounts of public data fine-tuning in Table 5. The result shows that at the same training round, there is no significant correlation between the public data ratio and the privacy-preserving capability. In addition, after 10 rounds of FedDIT, the base-data-only fine-tuning’s average ROUGE-L score all exceeds the instruction-augmented settings in four domains. Specifically, base-data-only fine-tuning exceeds augmented settings by 6.69%, 5.55%, 5.55%, and 6.16% on average, relatively.
The trend of the ROUGE-L score of base-data-only and augmented fine-tuning of each domain per round is shown in Figure 3. In specific, the augmented fine-tuning uses 5000 public data. We can see that the ROUGE-L score of augmented settings in each domain firstly increases and then starts decreasing or converging to a certain range. The base-data-only setting keeps the rising trend, especially the code domain, which indicates that as the number of training rounds increases, the base-data-only fine-tuning will remember more privacy information, while the risk of privacy leakage of the augmented fine-tuning will gradually decrease or converge.
6 Conclusion
We reveal that for FedDIT, the model performance is not linearly correlated with data heterogeneity, but is significantly impacted by cross-client domain coverage. To address this, we propose FedDCA, a domain-specific instruction augmentation method that maximizes domain coverage through greedy client center selection and retrieval-based augmentation. Additionally, FedDCA uses a heterogeneous encoder to reduce client-side computation overhead. Experiments across four domains demonstrate the effectiveness of FedDCA and the efficiency of FedDCA. Further analysis shows that as fine-tuning progresses, the risk of privacy leakage from augmented data decreases or converges.
Appendix A Appendix
A.1 Related Work
Federated Instruction Tuning. Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM (Zhang et al. (2024b); Wei et al. (2022)). Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios. So far, several FedIT frameworks(Ye et al. (2024b; a)) have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods. While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigate the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
PrivateLoRA(Wang et al. (2023b)) addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data. While FewFedPIT(Zhang et al. (2024d)) focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
Domain Instruction Augmentation. In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code(Nijkamp et al. (2023); Luo et al. (2023b)), medical(2024-Zhang-AlpaCare-), financial(Yang et al. (2023b; a); Zhang et al. (2023a); Wu et al. (2023)), mathematical(Yue et al. (2024); Luo et al. (2023a))). Existing works tend to use open-source domain-specific instruction tuning datasets for training. However, the target domain may not always have corresponding ready-made domain-specific instruction datasets. Even if they exist, these datasets are often limited in scale.
Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets(Wang et al. (2023a); Zhang et al. (2023c); Jiao et al. (2023); Lee et al. (2024); Xia et al. (2024)). For instance, Parrot(Jiao et al. (2023)) enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets. Furthermore, works like INSTA(Lee et al. (2024)) and LESS(Xia et al. (2024)) attempt efficient domain-specific instruction augmentation via dense retrieval. INSTA(Lee et al. (2024)) uses instructions without responses for effective retrieval. LESS(Xia et al. (2024)) assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval. 2) Using seed tasks for self-instruct(Luo et al. (2023b); Wan et al. (2023)): Explore-Instruct(Wan et al. (2023)), for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct. 3) Scaling instructions from the web: Recent works(Yue et al. (2024); Zhou et al. (2024)) highlight the immense potential of mining naturally occurring instructions from the internet. Compared to generated data, web-mined instructions exhibit less bias and greater diversity. MAmmoTH2(Yue et al. (2024)) firstly retrieves domain-relevant texts and employs a LLM to extract Q-A pairs, further refining them into instruction-response pairs. Jiuzhang3.0(Zhou et al. (2024)) distills GPT’s instruction generation capabilities into a smaller model and then uses it to generate instructions from the internet. In conclusion, models fine-tuned with augmented instructions have shown promising domain-specific capabilities.
Inspired by INSTA(Lee et al. (2024)), this paper uses a mixed dataset composed of code, medical, financial, mathematical, and general datasets to represent public data and does domain-specific instruction augmentation based on the client’s local instructions.
A.2 What Truly Counts in FedDIT
This section will first analyze the correlation between different non-iid levels and the model’s performance with separate experiments for both single and multiple domains in Section A.2.1. Further, to demonstrate the impact of non-iid on FedDIT with different test sets, we compare the performance of the global model trained on augmented data based on iid and non-iid cross-client data distribution. Additionally, we suggest that domain coverage is a key factor for FedDIT in Section A.2.2.
A.2.1 Data Heterogeneity Is Not Matter In FedDIT
Following the traditional approach of constructing different degrees of non-iid which are widely used in federated learning(Wang et al. (2020); Yurochkin et al. (2019)), we adopt Dirichlet distribution to construct various heterogeneity and use k-means with the cluster num $\xi=100$ to pseudo labeling instructions. Dirichlet distribution is affected by the hyperparameter $\alpha$, which enhances heterogeneity with a smaller $\alpha$ and decreases heterogeneity with a larger $\alpha$. We choose four widely used heterogeneity, which are $\alpha=[0.01,0.1,1,10]$(Ye et al. (2023); Zhang et al. (2023b); li2021model). Figure LABEL:fig:data_heterogeneity shows a visualization of the data distribution with different heterogeneity on the code dataset.
We perform instruction tuning both on single-domain and multi-domain with different heterogeneity. Specifically, for the single domain, we only use the in-domain data and then perform FedIT, and for multi-domain, we use the mixed dataset which consists of four domain-specific instruction datasets and a general instruction dataset, which are described in Appendix A.5. The training details are shown in Appendix A.6. Results for single domain and multi-domain are shown in Table 6. Both single-domain and multi-domain clearly show that the performance of LLM does not decrease due to the increase of data heterogeneity, but shows a nonlinear correlation, which indicates that the performance of LLM does not directly depend on data homogeneity, and exists other factors that play a key role.
Table 6: Performance(%) of different heterogeneity on single domain and multi-domain.
$\alpha$ Single Domain Multi-Domain
10 1 0.1 0.01 10 1 0.1 0.01
H-Eval 32.92 34.14 31.09 34.75 34.14 31.70 34.75 34.75
M-Med 70.6 69.9 71.0 71.3 71.7 71.3 69.9 71.3
FPB 67.90 72.11 73.76 68.89 60.06 61.63 60.97 62.04
FiQA 43.63 45.45 41.81 32.00 10.90 12.72 17.81 9.09
TFNS 74.37 76.00 75.71 72.69 65.36 65.24 64.69 64.61
GSM8K 53.75 51.40 51.93 52.76 46.5 43.5 41.5 45.5
Avg. 57.20 58.17 57.55 55.40 48.11 47.68 48.27 47.88
A.2.2 Domain Coverage: A Key Factor in FedDIT
Explore-Instruct(Wan et al. (2023)) enhances the coverage of domain-specific seed tasks through active exploration, then uses the self-instruct method for instruction data augmentation. This approach highlights the impact of domain coverage on domain-specific instruction tuning. Inspired by Explore-Instruct, we attempt to conduct more in-depth and extensive experiments to study the effect of domain coverage on FedDIT. Firstly we define the domain coverage of cross-client data in the FL setting. Assume the dataset of in-domain data $D^{d}$ represents the latent data distribution of this domain and the cross-client data is defined as $D^{c}=\cup_{k=1}^{N}\left(D_{k}^{l}\cup D_{k}^{g}\right)$. Inspired by the facility location function(Cornuejols et al. (1983)), we define the domain coverage of $D^{c}$ respect to $D^{d}$ as follows:
$d(D^{d},D^{c})=\frac{1}{\left|D^{c}\right|}\sum_{d\in D^{d}}\text{max}_{v\in(D^{c}\cap D^{d})}sim(d,v),$ (5)
where $sim(d,v)$ is the similarity between $d$ and $v$. Specifically, we use the cosine similarity function in FedDCA. Note that we only use the in-domain data in $D^{c}$ to calculate the domain coverage, because the out-of-domain data would mislead the domain coverage evaluation, as shown in Figure 7.
Figure 7: Misleadning of out-of-domain data on domain coverage calculation.
To better align with the real-world scenarios, we explore instruction augmentation based on both iid and non-iid cross-client data distribution and adopt direct data retrieval as described in Appendix A.3 for FedDIT. We set the number of clients to 10, while each client has 100 local instructions and obtains 5000 augmented public instructions from the server.
To construct iid data distribution, we randomly sample 1,000 from multi-domain datasets (code, medical, financial, mathematical, and general), which is detailed in Table 8 and divide them into 10 shards as each client’s local data. To construct non-iid data distribution, we perform k-means clustering with $\xi=100$. Each client randomly samples 100 instructions from different randomly selected clusters. Furthermore, for both iid and non-iid settings, direct data retrieval is performed based on the client’s local data.
Table 7: Performance(%) and domain coverage of iid and non-iid settings on different domains. The higher domain coverage correlates with better performance.
Test Set Performance (%) Domain Coverage
iid non-iid iid non-iid
H-Eval 35.36 33.53 0.8538 0.7994
M-Med 70.20 71.00 0.7800 0.8027
FPB 58.58 64.19 0.8523 0.9327
FiQA 17.09 19.27
TFNS 66.16 69.09
GSM8K 40.50 38.50 0.9137 0.8448
Table 7 presents the performance of FedDIT on different domains with iid and non-iid settings and shows the domain coverage in four domains. We can observe that both iid and non-iid settings outperform in some domains, but both collectively indicate that higher domain coverage correlates with better performance.
A.3 Direct domain data retrieval
In this section, we first show the detail of instruction-based dense retrieval in Appendix A.3.1, which is both used in direct retrieval and FedDCA. Then we explain the direct retrieval algorithm in Appendix A.3.2.
A.3.1 Instruction Based Dense Retrieval
[…]
$\textbf{Score}(\mathcal{E},I)=sim(\mathcal{E},w_{enc}(I)),$ (6)
where $sim(\cdot,\cdot)$ is the cosine similarity function. Based on the computed similarity between $\mathcal{E}$ and each $I$ in the public data, we then select the top-$N_{k}^{p}$ instructions as the retrieved public data for domain data augmentation.
A.3.2 Direct Retrieval
The direct domain data retrieval only utilizes instructions of the client’s local data without responses to perform the retrieval-based domain data augmentation. The detailed algorithm is described in Algorithm 2. In specific, for each client, we first encode the local instructions by the encoder model $w_{enc}$. Then we use the k-means algorithm to cluster the embeddings into $\xi$ clusters. Then the $\xi$ cluster centers are sent to the server and do the retrieval-based domain data augmentation. Then the retrieved public data is sent to the client for instruction tuning.
Algorithm 2 Direct domain data retrieval
1:Parameters:
2:Clients’ local datasets $D=\{D_{1},D_{2},\ldots,D_{N}\}$; Public dataset $D^{p}$; Local datasets $D^{l}=\{D^{l}_{1},D^{l}_{2},\ldots,D^{l}_{N}\}$; Number of clusters $\xi$; Encoder model $w_{enc}$; Pre-encoded public instruction embeddings $\mathcal{E}$.
3:$D^{g}\leftarrow\emptyset$
4:for $i\in\{1,2,\ldots,N\}$ do
5:     $\mathcal{E^{\prime}}\leftarrow\{w_{enc}(x)\mid x\in D_{i,instruction}^{l}\}$triangleright Encode the local instructions
6:     $\mathcal{C}\leftarrow\text{k-means}(\xi).fit(\mathcal{E^{\prime}})$triangleright Cluster local instructions, return cluster centers $\mathcal{C}$
7:     $\mathcal{S}\leftarrow\mathcal{C}\cdot\mathcal{E}^{T}$triangleright Compute the similarity score between $\mathcal{C}$ and $\mathcal{E}$
8:     $\mathcal{I}\leftarrow\emptyset$
9:     for $j\in\{1,2,\ldots,\xi\}$ do
10:         $\mathcal{S}^{\prime}\leftarrow\{s\mid s\in S_{j},\text{index}(s)\notin\mathcal{I}\}$triangleright Filter the selected indices
11:         $\mathcal{I^{\prime}}\leftarrow\text{indices of the top-}\frac{N_{k}^{p}}{\xi}\text{ elements in }\mathcal{S}^{\prime}$triangleright Retrieve the top $\frac{N_{k}^{p}}{\xi}$ indices
12:         $\mathcal{I}\leftarrow\mathcal{I}\cup\mathcal{I^{\prime}}$triangleright Update the selected indices
13:     end for
14:     $D^{g}_{i}\leftarrow\{D^{p}_{j}\mid j\in\mathcal{I}\}$triangleright Selected public instructions
15:     $D_{i}\leftarrow D^{g}_{i}\cup D_{i}^{l}$triangleright Obtain the augmented instruction dataset $D_{i}$
16:end for
A.4 Projector
To reduce the computational overhead of clients, we propose a heterogeneous encoder method called FedDCA. Inspired by LLaRA(Liao et al. (2023)), we utilize a projector to do the dimension alignment between the small encoder on the client side and the large decoder on the server side. As shown in Figure 8, where $d_{l}$ is the dimension of the client-side encoder, $d_{g}$ is the dimension of the server-side encoder.
Figure 8: The projector used on the server side for feature alignment, which consists of two fully connected layers and a ReLU activation layer in between.
A.5 Train and Test Dataset Information
Here we provide the train and test dataset details of code (CodeAlpaca [^2] https://huggingface.co/datasets/sahil2801/CodeAlpaca-20k), medical (MedAlpaca [^3] https://huggingface.co/datasets/medalpaca), financial (FinGPT(Yang et al. (2023a))) and mathematical (MathInstruct [^4] https://huggingface.co/datasets/TIGER-Lab/MathInstruct) domain, respectively.
Table 8: Train set information of each domain.
Dataset name Domain $N_{sample}$
CodeAlpaca Code 20,022
MedAlpaca Medical 33,955
FinGPT Financial 76,772
MathInstruct Mathematical 224,567
Alpaca General 52,002
As shown in Table 8, the public data consists of four domain-specific instruction datasets and a general instruction dataset, which are CodeAlpaca, MedAlpaca, FinGPT, MathInstruct, and Alpaca, respectively. For each domain’s FedDIT, we randomly select 1000 samples from the in-domain instruction and split them into 10 shards as each client’s local dataset. The rest instructions are used as the public dataset $D^{p}$.
Table 9: Test set information of each domain.
Dataset name Domain $N_{sample}$ Metric
HumanEval Code 164 Pass@1
MMLU-Med Medical 1,089 Acc
FPB Financial 152 Acc
FiQA Financial 35 Acc
TFNS Financial 299 Acc
GSM8K Mathematical 1,319 Exact Match
Table 9 shows the detail of the test set used for evaluation, HumanEval is used for the code domain, MMLU-Med is used for the medical domain, FPB, FiQA and TFNS are used for the financial domain, and GSM8K is used for the mathematical domain. Specifically, MMLU-Med uses subjects anatomy, clinical_knowledge, college_biology, college_medicine, medical_genetics and professional_medicine in MMLU.
A.6 Implementation Details
We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round. Then each available client performs FedDIT for 10 steps with AdamW optimizer and the batch size $B=32$ in a round. The initial learning rate is $55e-$ 5 with a cosine learning rate scheduler. Our experiment utilizes the widely used LLM, Llama3-8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B as the base model with 2048 max sequence length and adopts LoRA tuning method. The rank of LoRA is 16 and the scalar alpha is 16. For k-means(Vardakas & Likas (2023)), we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7. For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning(Khosla et al. (2020)). We utilize bge-large-en-v1.5666https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions. While we utilize all-MiniLM-L6-v2777https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 as the client’s small encoder which outputs embeddings of 384 dimensions and bge-large-en-v1.5 as the server’s encoder for FedDCA.
A.7 Prompts used in the Self-Instruct data generation
To generate the Self-Instruct data, we prompt GPT-3.5 to generate the instruction with the designed prompt in Figure 9. Specifically, we randomly sample two examples from the client’s local data to guide GPT-3.5 generating the in-domain instruction, and sample one example from the client’s local data for one-shot in-context learning to guide GPT-3.5 generating responses into the example’s format.
Figure 9: Prompts used in the Self-Instruct data generation. (a) Prompt for generating new instructions. Two examples are randomly sampled from the client’s local data for in-context demonstration. (b) Prompt for generating responses. We prompt GPT-3.5 to generate responses with a randomly selected example for one-shot in-context learning.
A.8 Prompt used for memory extraction attack
As we use Llama3-8B as our base model and format the instructions and responses into the Alpaca’s format, to utilize the auto-regression nature of LLM to extract the instruction, we prompt the model to generate the instruction using the prompt in Figure 10, which is exactly the prefix of the Alpaca’s template.
Figure 10: Prompt used for memory extraction attack.
A.9 Further Analysis
To further study the effect of different hyperparameters of FedDCA, we undertake a thorough analysis including various retrieval amounts, and different k-means cluster numbers $\xi$. In addition, we perform the ablation study on whether using the similarity threshold $\alpha$ in the greedy client center selection.
Effect of Retrieval Number. We report the performance and the corresponding domain coverage of FedDCA with different retrieval amounts on the four domains in Figure 11, respectively. We can see that the domain coverage of FedDCA is increasing along with the retrieval amount in different trends, as well as the performance. Specifically, the domain coverage of each domain increases by 6.36%, 18.69%, 5.04%, and 8.06% in relative, respectively. Along with domain coverage increasing, the performance of FedDCA is increasing by 3.05%, 2.20%, 4.08%, and 6.95% for each domain. Noted that although the code domain is more about following a certain paradigm, which could perform well with a few data and more fine-tuning rounds, it still could benefit from the instruction augmentation.
Figure 11: Effect of different retrieval amounts on the performance of FedDCA and its domain coverage. We show the results on four domains separately. Here we use the FPB test set to evaluate the performance in the financial domain.
Impact of Different Cluster Number. The hyperparameter $\xi$ is the number of clusters in the k-means algorithm. The experiment is conducted on $\xi=[N,2N,4N,8N]$, where $N$ is the number of clients. Following the setting in Appendix A.6, we set $N=10$. We report the domain coverage of the augmented dataset via FedDCA with different cluster numbers $\xi$ on the four domains in Figure 12, respectively. Results show that there is no best $\xi$ for all domains. Specifically, the best $\xi$ of code, medical, financial, and mathematical domains are 80, 80, 40, and 10, respectively.
Figure 12: Impact of different cluster number $\xi$ on the cross-client domain coverage. We report the domain coverage of the augmented dataset via FedDCA with different cluster numbers $\xi$ on the four domains.
Ablation Study. We conduct the experiment with FedDCA w/o similarity threshold $\alpha$ on the four domains based on the FedAvg FL strategy. The performance and the corresponding domain coverage are shown in Table 10, where FedDCA without the similarity threshold $\alpha$ is marked as FedDCA. The result shows that the performance of FedDCA with similarity threshold $\alpha$ is slightly better than FedDCA without using $\alpha$ in code, financial, and mathematical domains, as the similarity scores in the medical domain are relatively lower. We show the similarity score distribution of the four domains in Figure LABEL:fig:similarity_score. For each domain, we plot each similarity score’s distribution of 10 clients. The similarity score is computed between the the selected client center and the public data, then we show the similarity score distribution using the histogram plot.
Table 10: Ablation study on the performance and domain coverage of FedDCA w/o similarity threshold $\alpha$.
Metric Performance(%) Domain Coverage
FedDCA FedDCA FedDCA FedDCA
H-Eval 35.97 36.58 0.8972 0.9348
M-Med 73.40 73.40 0.9348 0.9348
FPB 66.25 67.24 0.9353 0.9815
FiQA 23.27 35.27
TFNS 69.34 73.32
GSM8K 51.78 52.46 0.9128 0.9320
A.10 Augmentation strategy visualization
To more intuitively compare the domain coverage of different instruction augmentation methods, we randomly sampled 5,000 instructions obtained through these methods and randomly sampled 10,000 in-domain instructions as the background, which represents the distribution of specific domains in the public dataset. We then visualized the results using t-SNE(van der Maaten & Hinton (2008)), as shown in Figure LABEL:fig:augmentation_strategy_visualization. The plot shows that FedDCA encompasses the most of in-domain data, which is consistent with FedDCA’s domain coverage of each domain shown in Table 3. Also, we can observe that the random sampling strategy selects a lot of out-of-domain data while does not have good coverage in specific domains.
A.11 Frequently Used Notation
Table 11: Frequently used notation.
NOTATIONS REMARK
$N,C$ Clients number, client set $C=\{c_{1},c_{2},\ldots,c_{N}\}$.
$D^{p},D^{d},D^{c},D^{l}$ The public datasets, the in-domain data, the cross-client dataset, client’s local private data.
$N_{k}^{l},N_{k}^{p}$ Number of local private data on the $k$-th client, number of the retrieved public data on the $k$-th client.
$D_{k}^{l},D_{k}^{g},D_{k}$ The local private data on the $k$-th client, the retrieved public data on the $k$-th client, the augmented dataset on the $k$-th client.
$\Lambda,\mathcal{P},\mathcal{C}$ A specific sampling strategy that performs instruction augmentation on the server side, selected client center set, the cluster centers obtained locally and sent to the server for the greedy client center selection.
$\phi,\Delta\phi$ LLM’s pre-trained parameters, additional LoRA parameters.
$w,w_{enc},w_{p}$ Merged model parameters from the frozen LLM’s parameters $\phi$ and the additional LoRA parameters $\Delta\phi$, encoder model, projector model.
$F_{k}(w;\mathcal{D}),l(w;x,y)$ Loss of model $w$ over a specific dataset $\mathcal{D}$, the instructon tuning loss of model $w$ over a data sample $\left(x,y\right)$.