### v5

# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

Abstract
Federated Domain-specific Instruction Tuning (FedDIT) utilizes limited cross-client private data together with various strategies of instruction augmentation, ultimately boosting model performance within specific domains. To date, the factors affecting FedDIT remain unclear, and existing instruction augmentation methods primarily focus on the centralized setting without considering distributed environments. Our experiments reveal that the cross-client domain coverage, rather than data heterogeneity, drives model performance in FedDIT. In response, we propose FedDCA, which optimizes domain coverage through greedy client center selection and retrieval-based augmentation. At its core, the greedy selection procedure iteratively picks client centers that maximize the diversity and coverage of the instruction space while avoiding redundancy with previously selected centers. This ensures broad yet efficient coverage of the domain distribution across clients. For client-side computational efficiency and system scalability, FedDCA, the variant of FedDCA, utilizes heterogeneous encoders with server-side feature alignment. Extensive experiments across code, medical, financial, and mathematical domains substantiate the effectiveness of both methods, as well as plug-and-play capability. We further analyze privacy preservation against memory extraction attacks, showing that while privacy leakage risk is independent of augmented public data ratio, it decreases or converges as training progresses.
1 Introduction
Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to facilitate collaborative training of large language models (LLM) in distributed environments while maintaining the confidentiality of private data McMahan et al. (2017); Ye et al. (2024b); Zhang et al. (2023b). This methodology allows for the exchange of model parameters among distributed data holders, thereby achieving a careful balance between privacy preservation and efficient model optimization. Despite the establishment of various FedIT frameworks Ye et al. (2024b); Kuang et al. (2023); Zhang et al. (2023b), existing literature has not adequately addressed the practical challenges that Federated Domain-specific Instruction Tuning (FedDIT) may encounter in real-world applications. For instance, FedIT generally necessitates a sufficient amount of instruction data for fine-tuning, which is often a shortage in domain-specific fine-tuning contexts Zhang et al. (2024b).
Figure 1: Visualization of different instruction augmentation strategies in the mathematical domain. The Random Sampling strategy randomly samples instructions from the public dataset. Direct Retrieval performs dense retrieval from the public dataset based on the client’s local data distribution without considering the cross-client domain coverage. FedDCA strategically selects client centers to maximize cross-client domain coverage and then performs domain data retrieval. The t-SNE visualization shows FedDCA has a better domain coverage than other sampling strategies.
In this study, we investigate FedDIT, a novel approach within the FL paradigm aimed at boosting the performance of LLMs in specific domains. Unlike general FedIT, which enhances model effectiveness across diverse tasks without accounting for local data shortages, FedDIT addresses the unique challenge of clients only having limited domain-specific data. To overcome this, FedDIT enriches local data through a specific instruction augmentation strategy, crucial for effective instruction tuning and avoiding performance degradation. Our results reveal that, except for the code domain, where a standardized paradigm is primarily adhered to, relying solely on high-quality in-domain local data can be insufficient due to limited scale, leading to a decline in performance (see underlined values in Table 2). FedDIT aims to develop a domain-specific LLM through collaborative training and instruction augmentation while safeguarding client privacy, ensuring proficiency in tasks pertinent to its designated domain. For simplification and better instruction quality Zhang et al. (2024b); Toshniwal et al. (2024), we focus on a specific scenario of FedDIT involving a server-hosted public dataset, abstracting open-source instruction datasets on the web across multiple domains. This dataset is used for various instruction augmentation strategies, enhancing model performance in specific domains.
Additionally, the factors affecting FedDIT are still unclear. Compounding this uncertainty, introducing augmented instructions may further complicate results, making it difficult to ascertain effective improvement strategies. We conduct experiments to unveil a significant finding: there is no monotonic correlation between the degree of non-independent and identically distributed (non-iid) and LLM’s performance in the context of FedDIT. Inspired by Explore-Instruct Wan et al. (2023), which highlights the potential of domain coverage in domain-specific instruction tuning. Unlike previous metric that measures domain coverage using the distribution of verb-noun pairs, we define a more general, representation-based cross-client domain coverage metric and investigate its impact on FedDIT. Our results show that domain coverage significantly influences model performance within the corresponding domain.
To maximize the cross-client domain coverage without compromising client data privacy, we propose a novel FedDIT algorithm, Federated Instruction Tuning of LLMs with Domain Coverage Augmentation, termed FedDCA. This algorithm employs a greedy client center selection process and implements instruction augmentation through dense retrieval on the server side. The fundamental idea of FedDCA is to select client centers to expand the diversity and coverage of augmented instruction datasets within a specific domain (as shown in Figure 1). By strategically optimizing domain coverage at each step, FedDCA efficiently constructs the augmented train set that enhances both the learning and generalization capabilities of the model, leading to superior performance on domain-specific tasks. Furthermore, to mitigate computational overhead on the client side and enhance the system scalability, we propose FedDCA, which employs heterogeneous encoders of different sizes and capacities. To achieve feature alignment, we train a projector on the server side using public data and employ contrastive learning techniques.
We demonstrate the effectiveness and plug-and-play capability of FedDCA through comprehensive experiments conducted across four domains: code, medical, financial, and mathematical. These are compared against a range of baselines, which can be categorized into unaugmented and augmented methods. Results show that FedDCA outperforms other methods from at least 0.84% to the maximum of 29.19% in terms of model performance, with a relative domain coverage improvement from 4.82% to 21.36%. Extensive experiments across privacy analysis, computational efficiency and other further investigations are also conducted.
2 Problem Formulation
Federated Domain-specific Instruction Tuning (FedDIT) is a federated learning approach designed to improve the performance of LLMs in specific domains by utilizing limited cross-client private data in combination with domain-specific instruction augmentation strategies Zhang et al. (2024b); Xia et al. (2024); Wang et al. (2022). We abstract the public data as a server-hosted multi-domain dataset (the composition is detailed in Table 1). The presence of a public dataset on the server is only one possible scenario, and the proposed algorithm is independent of the presence of a public dataset on the server. Even if the server does not have a public dataset, clients can retrieve public instructions by themselves based on the received client centers from the server or perform self-instruct, thereby achieving data augmentation.
Consider $N$ distributed clients, each with local private data $D_{k}^{l}$ of size $N_{k}^{l}$, and augmented data $D_{k}^{g}$ from the server’s public dataset $D^{p}$. Due to constraints like memory and computation, client $c_{k}$ can accept at most $N_{k}^{p}$ public instructions. The server maintains $D^{p}$ that spans multiple domains and is responsible for data augmentation strategies $\Lambda$ and parameter aggregation. For efficiency, we adopt Low-Rank Adaption (LoRA) Hu et al. (2022) as the fine-tuning method, tuning additional parameters $\Delta\phi$ while keeping pre-trained LLM parameters $\phi$ frozen.
The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data Ye et al. (2024b); Zhang et al. (2024b, 2023b), which is defined as:
$\small\operatorname*{arg\,min}_{\Delta\phi}\left\{F(\phi,\Delta\phi)\triangleq\sum_{k=1}^{N}p_{k}F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{l},D_{k}^{g};\alpha_{k}\right)\right\},$ (1)
where $F_{k}(\cdot)$ is the accumulated instruction fine-tuning loss of model $P_{\phi+\Delta\phi_{k}}$ computed on the client $c_{k}$’s augmented dataset $D_{k}^{l}\cup D_{k}^{g}$. $p_{k}$ is client $k$’s weight based on data ratio, and $\alpha_{k}=\frac{N_{k}^{p}}{N_{k}^{l}+N_{k}^{p}}$ is the ratio of public data. Eq.1 minimizes the summed empirical loss across clients’ augmented instructions to pursue the in-domain utility of the obtained global model. The client $c_{k}$’s empirical loss $F_{k}(\phi,\Delta\phi_{k};\mathcal{D})$ is calculated as:
$F_{k}(\phi,\Delta\phi_{k};\mathcal{D})\triangleq\frac{1}{|\mathcal{D}|}\sum_{j=1}^{|\mathcal{D}|}l(P_{\phi+\Delta\phi_{k}};x_{j}),$ (2)
where $x_{j}\in D,\forall j\in\{1,2,\dots,|D|\}$. The instruction tuning loss $l(\cdot;\cdot)$ on a sample $(x,y)$ is defined as $-\sum_{t=1}^{\left|y\right|}\log\left(w\left(y_{t}|x,y_{<t}\right)\right)$, where $x$ is the formated instruction and $y$ is the corressponding response.
In summary, we formalize FedDIT as a novel FL framework that focuses specifically on domain-specific instruction tuning. In contrast, FedIT typically demands a substantial quantity of instructions for fine-tuning. Domain-specific instruction augmentation can be divided into two categories: 1) Generative methods, which generate instructions through the pre-trained LLM locally or external APIs. 2) Retrieval-based methods, which retrieve instructions from the web. The former methods either compromise privacy or have high computational overhead. Thus, in this work, we focus on retrieval-based methods to utilize diverse and high-quality public data.
3 What Truly Counts in FedDIT
(a) Exp1. Code
(b) Exp1. Med
(c) Exp1. Fin
(d) Exp1. Math
(e) Exp2. Performance (%)
(f) Exp2. Domain Coverage
Figure 2: (a)-(d) Performance (%) of different heterogeneity in each domain with 10 and 100 clients. (e)-(f) Performance and domain coverage of iid and non-iid settings on different domains. Experiments show that cross-client domain coverage significantly impacts model effectiveness, while data heterogeneity shows no monotonic correlation with performance.
Data Heterogeneity Is Not Matter in FedDIT.
We adopt the Dirichlet distribution Wang et al. (2020); Yurochkin et al. (2019) to construct various heterogeneity ($\alpha=[0.01,0.1,1,10]$) and use k-means with $\xi=100$ to pseudo labeling instructions.
We perform instruction tuning on both 10 and 100 clients with 2 randomly selected clients participating in each round. For each domain, we only use the in-domain data and then perform FedIT. We perform the experiments 3 times with different random seeds (42, 43 and 44) and report the average performance and the standard deviation in each domain. As shown in Figures 2(a), 2(b), 2(c) and 2(d), the performance of LLM does not decrease due to the increase of data heterogeneity but shows a non-monotonic correlation, which indicates that the performance of LLM does not directly depend on data heterogeneity and other factors that play a key role.
Domain Coverage: A Key Factor in FedDIT.
Different from Explore-Instruct, which defines domain coverage through the distribution of verb-noun pairs, we attempt to conduct more in-depth and extensive experiments to study the effect of domain coverage on FedDIT. Firstly, we define the domain coverage in the FL setting, considering the cross-client data distribution. Assume the dataset of in-domain data $D^{d}$ represents the latent data distribution of this domain and the cross-client data is defined as $D^{c}=\cup_{k=1}^{N}\left(D_{k}^{l}\cup D_{k}^{g}\right)$.
Inspired by the facility location function Cornuejols et al. (1983), we define the domain coverage of $D^{c}$ respect to $D^{d}$ as follows:
$\small d(D^{d},D^{c})=\frac{1}{\left|D^{d}\right|}\sum_{d\in D^{d}}\text{max}_{v\in(D^{c}\cap D^{d})}sim(d,v),$ (3)
where $sim(d,v)$ is the cosine similarity between $d$ and $v$. Note that we only use the in-domain data in $D^{c}$ for the domain coverage calculation to avoid misleading.
To better align with the real-world scenarios, we explore instruction augmentation based on both iid and non-iid cross-client data distribution and then perform dense retrieval for FedDIT. We set the number of clients to 10, while each client has 100 local instructions and obtains 1000 augmented public instructions from the server. For the iid setting, we randomly sample 1k from the public dataset and divide them into 10 shards as each client’s local data. For the non-iid distribution, we perform k-means clustering with $\xi=100$. Each client randomly samples 100 instructions from randomly selected distinct clusters.
Figures 2(e) and 2(f) presents the performance and the according domain coverage of FedDIT in different domains with iid and non-iid settings. We can observe that both iid and non-iid settings outperform in some domains, but both collectively indicate that higher domain coverage correlates with better performance.
4 Method
Figure 3: Overview of FedDCA, which consists of three stages: 1) The client $c_{k}$ performs local instructions clustering and sends cluster centers $\mathcal{C}_{k}$ to the server. 2) The server first does greedy client center selection to maximize domain coverage (obtains $\mathcal{P}$) and performs client-center-based domain data retrieval, then sends the augmented instructions $D_{k}^{g}$ to the client $c_{k}$. 3) Clients fine-tune the LLM collaboratively, and the LoRA parameters $\Delta\phi$ are exchanged between clients and the server.
Based on the above empirical observations, we propose FedDCA, which enhances domain coverage to obtain a LLM that performs well on domain-specific tasks (shown in Figure 3). We first formulate and analyze the cross-client domain coverage optimization problem. Then, we introduce the FedDCA algorithm, which consists of two main stages: greedy client center selection and client-center-based domain data retrieval. For computational efficiency and system scalability, we introduce a variant of FedDCA with heterogeneous encoder settings, named FedDCA.
4.1 Optimization Problem
As domain coverage directly affects the in-domain performance of the LLM, FedDCA aims to maximize the domain coverage of the cross-client augmented data $D^{c}$ with respect to the in-domain data distribution $D^{d}$. However, as clients can not send the local data to the server for privacy, directly finding a cross-client dataset that maximizes the domain coverage is unrealistic. Instead, finding a proper client center set $\mathcal{P}$ with optimized domain coverage and using it to perform instruction retrieval on the public data is an approximation problem.
Suppose there exists a metric space $\mathcal{X}$, a set of cross-client local instruction embeddings $\mathcal{E}\in\mathcal{X}$, and a set of cluster centers $\mathcal{P}\in\mathcal{X}$. Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm Wu (2012). Additionally, in the federated setting, communication cost is always a critical factor. Thus, we formulate the optimization problem as follows:
$\footnotesize\operatorname*{arg\,min}_{\mathcal{P}}$$\left\{\sum_{i=1}^{N}|\mathcal{P}_{i}|-\frac{1}{|D^{d}|}\sum_{d\in D^{d}}\left(\max_{p\in\mathcal{P}}\operatorname{sim}(d,p)\right)\right\},$ (4)
s.t. $\quad|\mathcal{P}_{i}|\leq\xi,\quad\forall i\in\{1,2,\dots,N\}$ (5)
The second term is the domain coverage of the selected client center set $\mathcal{P}$, and $sim(\cdot,\cdot)$ is the cosine similarity function. However, this optimization problem is NP-hard.
For simplicity, we let $\xi$ be the constant $N$ (further discussed in Section 6.3). To solve this problem, the time complexity is $\mathcal{O}\left(C_{\xi N}^{N}\left(\mathcal{N}\left(N+N\log N\right)\right)\right)$, where $\mathcal{N}$ is the size of the public data $D^{p}$. As it is a factorial equation, the computational cost increases exponentially with $N$. Furthermore, the enormous amount of public data results in a large $\mathcal{N}$. To enhance computational efficiency, we propose a greedy algorithm as described below.
4.2 FedDCA
As the goal of FedDCA is to find a proper client center set $\mathcal{P}$ to optimize the objective defined in Eq. 4, we propose a greedy algorithm as shown in Algorithm 1 to solve this problem in polynomial time and obtain a sub-optimal solution. Given the cluster centers $\{\mathcal{C}_{i}\}_{i=1}^{N}$ received from each client, which are obtained by clustering local instructions. FedDCA on the server consists of two main steps: greedy client center selection and client-center-based domain data retrieval.
Naive Greedy Client Center Selection.
In this paper’s FedDIT setting, a natural attempt to achieve greedy domain coverage optimization is using the public dataset $D^{p}$ on the server and maximizing $d(D^{d},\mathcal{P})$. Here, each client provides a set of candidate embeddings, and the server greedily picks (or updates) the cluster center to maximize domain coverage on $D^{p}$. While this approach is intuitive, it has three main limitations: 1) High computation complexity when $D^{p}$ is large. Each step of finding the best center $q^{*}$ (to add or update) must re-evaluate coverage w.r.t. $|D^{d}|$ data points, often incurring complexity on the order of $\mathcal{O}\bigl{(}N\xi|D^{d}|(N+N\log N)\bigr{)}$ per selection step. This quickly becomes prohibitively expensive for large $|D^{d}|$. 2) Difficulty in domain coverage computation. The server may have limited knowledge about which domain each public instruction belongs to, or only know it partially. Additionally, the definition of domain is flexible without unified annotation standards - clients’ perceptions of domains may differ from the server’s understanding. This makes it challenging for the server to accurately compute domain coverage. 3) No guaranteed public data. In many realistic FedDIT deployments, the server may not actually have an in-domain public dataset. Clients might rely on external retrieval or self-instruct generation for data augmentation, leaving no $D^{p}$ accessible on the server. The coverage-based greedy approach is then infeasible.
Cluster-center-based Client Center Selection.
To address these issues, we propose an alternative method that does not rely on public dataset. We denote the full set of these uploaded cluster centers $\{\mathcal{C}_{i}\}_{i=1}^{N}$ as $\mathcal{C}_{\text{all}}$, which is the proxy for domain coverage space. Since each client must pick one cluster center from $\mathcal{C}_{\text{all}}$. We then define the coverage score for client center selection:
$\quad C(\mathcal{P})$ $=\sum_{z\,\in\,\mathcal{C}_{\text{all}}}\max_{p\in\mathcal{P}}\mathrm{sim}(z,p),$ (6)
$where\quad\mathcal{P}$ =$=\{\,p_{1},\dots,p_{N}\,\},\quad p_{i}\in\mathcal{C}_{\text{all}},$ (7)
and seek to optimize it. This avoids relying on $D^{d}$, greatly reducing coverage-computation overhead and allowing operation without a public dataset.
Specifically, each client $c_{i}$ locally runs $k$-means to produce $\xi$ candidate centers $\mathcal{C}_{i}$, then uploads these centers to the server. The server initializes a selection $\mathcal{P}$ by picking one cluster center from each client (thus $|\mathcal{P}|=N$) and computes the overall coverage solely based on these uploaded centers. In each iteration, the algorithm employs a coordinate-ascent style update: it removes the currently chosen center from one client’s portion of $\mathcal{P}$, tries replacing it with other candidate cluster centers, and keeps the replacement if it improves the coverage; this process is repeated until no further improvement is found (as shown in Algorithm 1).
Algorithm 1 FedDCA: Greedy Client Center Selection
1:Initialize:
2: For each client $c_{i}$, pick any initial center $p_{i}\in\mathcal{C}_{i}$.
3: Let $\mathcal{P}=\{p_{1},\dots,p_{N}\}$ and compute its coverage $C(\mathcal{P})$.
4:$\Omega^{*}\leftarrow C(\mathcal{P})$
5:for $i=1,\dots,N$ do
6:    $\mathcal{P}_{-i}\leftarrow\mathcal{P}\setminus\{p_{i}\}$triangleright Remove $p_{i}$ from the current center set
7:    $\nu^{*}\leftarrow p_{i}$
8:    for $q\in(\mathcal{C}_{\text{all}}-\mathcal{P}_{-i})$ do
9:         $C_{i}=C(\mathcal{P}_{-i}\cup\{q\})$
10:         if $C_{i}>\Omega^{*}$ then
11:             $\Omega^{*}\leftarrow C_{i}$; $\nu^{*}\leftarrow q$
12:         end if
13:    end for
14:    $p_{i}\leftarrow\nu^{*}$; $\mathcal{P}\leftarrow\mathcal{P}_{-i}\cup\{p_{i}\}$; $C(\mathcal{P})\leftarrow\Omega^{*}$
15:end for
16:Output: The final center set $\mathcal{P}$ and its coverage $C(\mathcal{P})$.
Since coverage is computed over just $N^{2}$ local centers (i.e., $\mathcal{C}_{\text{all}}$) rather than a large $D^{p}$, the computational overhead is much more manageable. Moreover, by not relying on a public dataset, this approach is applicable for various scenarios, including both data augmentation and data selection.
Domain Data Retrieval.
For each client center $p_{k}$, the server performs dense retrieval (detailed in Appendix A.2) on public dataset $D^{p}$ to get the top-$N_{k}^{p}$ similar public instructions, then sends retrieved public datasets $\{D_{1}^{g},\dots,D_{N}^{g}\}$ to each clients. Specifically, to avoid the overlap between public data and local private data, we set a threshold $\alpha$ to filter the public instructions that have a similarity score larger than $\alpha$ with the client center.
FedDCA: Heterogeneous Encoders.
To enhance client-side computational efficiency and system scalability, we introduce FedDCA. This method employs a small encoder $\omega$ on the client side to reduce computational overhead and a larger encoder $\omega^{\prime}$ on the server side. However, the output dimension of heterogeneous encoders may not be consistent. Therefore, we introduce a projector $w_{p}$ on the server for feature alignment, which consists of two fully connected layers and a ReLU activation layer in between. We use contrastive learning Khosla et al. (2020) and train $w_{p}$ on the public instructions. For $\forall{x_{i}}\in D^{p}$, $\omega$ outputs embedding $h_{i}$, $\omega^{\prime}$ outputs embedding $h_{i}^{\prime}$ and $w_{p}$ outputs embedding $\varphi_{i}$. For input $x_{i}$, the positive sample pair is $(h_{i}^{\prime},\varphi_{i})$ and negative sample pairs are $\{\varphi_{j},j\neq i\}$. Let the batch size be $B$ and $\tau$ be the temperature parameter, then the training objective is defined as follows:
$\footnotesize\mathcal{L}=-\sum_{i=1}^{B}\log\frac{\exp\left(\text{sim}(\varphi_{i},h^{\prime}_{i})/\tau\right)}{\sum_{j=1}^{B}\mathbb{I}_{[j\neq i]}\exp\left(\text{sim}(\varphi_{i},h^{\prime}_{j})/\tau\right)},$ (8)
In summary, FedDCA leverages greedy client center selection and domain data retrieval to maximize cross-client domain coverage, leading to better model performance on domain-specific tasks. Additionally, FedDCA employs heterogeneous encoders with server-side feature alignment to reduce client-side computational overhead while maintaining effectiveness.
5 Discussions
Computation & Communication.
Through greedy client center selection, FedDCA solves the optimization problem defined in Eq. 4 in the polynomial time. As mentioned in Section 4.1, the time complexity of greedy client center selection is $\mathcal{O}\left(N^{4}\xi^{2}\left(1+\log N\right)\right)$. For domain data retrieval, its time complexity is $\mathcal{O}\left(N(\mathcal{N}\log\mathcal{N})\right)$. In addition, the heterogeneous encoder setting FedDCA further reduces the computation overhead on the client side.
In the domain instruction augmentation stage, each client sends $\xi$ cluster centers to the server. The server returns $N_{k}^{p}$ retrieved public instructions to the client $c_{k}$. FedDCA follows standard FedIT procedures, efficiently exchanging LoRA parameters $\Delta\phi_{k}$ between clients and the server.
Privacy.
Comparing FedDCA with other FedIT methods Zhang et al. (2024b); Ye et al. (2024b), the difference lies in the greedy client center selection. In this stage, the client only uploads the cluster center to the server, which is the average of embeddings to its cluster. In addition, the potential privacy leakage can be further avoided through homomorphic encryption Acar et al. (2018), which allows the server to directly compute on ciphertext for matrix multiplication for dense retrieval.
6 Experiments
To demonstrate the effectiveness of FedDCA and its variant FedDCA, we conduct extensive experiments across various domains and with several baselines. For additional results and analysis, please refer to Appendix C.
6.1 Experimental Setup
Alpaca CodeAlpaca HumanEval MedAlpaca MMLU-Med
$N_{sample}$ 52,002 20,022 164 33,955 1,089
Metric - - Pass@1 - Acc
FinGPT FPB FiQA TFNS MathInstruct GSM8K
$N_{sample}$ 76,772 152 35 299 224,567 1,319
Metric - Acc Acc Acc - Exact Match
Table 1: Dataset information of each domain. The test set is marked by $\ddagger$. The public dataset is composed of these five train sets. We denote HumanEval and MMLU-Med as H-Eval and M-Med for short.
Dataset and Evaluation Metrics.
As shown in Table 1, to evaluate the performance of FedDCA, we conduct experiments on four domains: code, medical, financial, and mathematical. We set the number of clients to 10 by default, while each client has 100 local instructions and obtains 5k augmented public instructions from the server.
Baselines.
We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented methods, including zero-shot inference and FedIT, which are composed of four widely used FL methods, including FedAvg McMahan et al. (2017), FedProx Li et al. (2020), SCAFFOLD Karimireddy et al. (2020) and FedAvgM Hsu et al. (2019). 2) Augmented methods, including Random Sampling, Direct Retrieval, LESS Xia et al. (2024), and Self-Instruct Wang et al. (2022). Specifically, Random Sampling randomly samples $N_{k}^{p}$ public data for each client, while Direct Retrieval is a simplified version of FedDCA that directly retrieves public data based on each client’s cluster center set without greedy client center selection, not considering the cross-client domain coverage.
6.2 Main Results
Performance.
Table 2 shows the performance and the corresponding domain coverage of FedDCA and other baselines on four domains. FedDCA outperforms the other nine baselines in all domains, with a substantial improvement from at least 0.84% to the maximum of 29.19% over other baselines. In particular, FedDCA+FedAvg and FedDCA+FedProx perform better across four domains. We see that although FedDCA has a performance drop than FedDCA, it still outperforms other baselines. We can observe that FedIT methods perform well in the code domain, even better than most augmented methods. This could be attributed to two reasons: 1) The code domain follows a certain paradigm. 2) With limited local data, unaugmented methods learn the same data more times given the same epoch and batch size, which may be an unfair comparison. Nonetheless, FedDCA still surpasses the best baseline in the code domain.
Domain Coverage.
FedDCA achieves the highest domain coverage across all domains, surpassing other baselines by 4.82% to 21.36% on average. Although FedDCA uses a smaller encoder on the client side for computational efficiency, it results in a slight 2.89% drop in average domain coverage but still outperforms the best baseline by 2.78% on average. These results demonstrate the effectiveness of both FedDCA and FedDCA in enhancing domain coverage.
Method H-Eval M-Med FPB FiQA TFNS GSM8K
Zero-shot 29.88/- 70.60/- 55.94/- 18.54/- 59.21/- 23.27/-
FedAvg McMahan et al. (2017) 39.03/0.8126 68.40/0.6990 58.25/0.8529 14.18/0.8529 66.62/0.8529 47.46/0.7871
FedProx Li et al. (2020) 37.20/- 69.10/- 56.51/- 14.90/- 66.45/- 47.15/-
SCAFFOLD Karimireddy et al. (2020) 37.80/- 70.20/- 62.71/- 15.27/- 66.49/- 49.27/-
FedAvgM Hsu et al. (2019) 32.32/- 64.70/- 68.14/- 29.27/- 70.32/- 46.85/-
Random Sampling 32.93/0.8512 71.30/0.7940 64.19/0.9196 13.09/0.9196 65.53/0.9196 47.38/0.8651
Direct Retrieval 34.14/0.9396 72.20/0.8830 66.31/0.9293 19.11/0.9293 67.62/0.9293 50.87/0.8967
LESS Xia et al. (2024) 28.04/0.8509 71.00/0.7737 60.56/0.8917 16.00/0.8917 61.14/0.8917 43.13/0.8352
Self-Instruct Wang et al. (2022) 32.92/0.8966 71.90/0.8586 59.73/0.9015 20.67/0.9015 66.54/0.9015 50.79/0.8811
FedDCA 34.75/0.9532 73.30/0.8972 67.10/0.9538 30.54/0.9538 71.01/0.9538 51.55/0.9096
FedDCA+FedAvg 36.58/0.9766 74.50/0.9348 67.24/0.9815 35.27/0.9815 73.32/0.9815 52.46/0.9320
FedDCA+FedProx 32.92/- 72.40/- 72.93/- 38.18/- 77.55/- 51.25/-
FedDCA+SCAFFOLD 39.87/- 73.20/- 72.68/- 33.09/- 75.50/- 50.26/-
FedDCA+FedAvgM 33.53/- 68.90/- 71.45/- 31.45/- 72.52/- 49.76/-
Table 2: Performance (%) and domain coverage of FedDCA, FedDCA and other nine baselines. Due to limited local data or inappropriate data augmentation strategy, occurs performance degradation in the Unaugmented and Random Sampling settings (underlined values). FedDCA outperforms all other baselines.
Plug-and-Play.
To demonstrate the plug-and-play capability of FedDCA with other augmentation methods, we first perform instruction sampling through different baselines to obtain augmented data (5k instructions per client) and then utilize FedDCA to conduct data selection (select 1k instructions per client) for FedDIT. As shown in Table 3, using FedDCA in combination with other augmentation methods still achieves comparable performance despite using only 20% of the sampled data. FedDCA efficiently selects the most representative samples based on domain coverage, enabling effective training with less data while maintaining performance.
Computation Analysis.
To show the computational efficiency of FedDCA, we compare these to FedDCA from two aspects: 1) Encoder model size. 2) Encoding time overhead. We test the encoders’ encoding time on the public dataset. As shown in Figure 6(a), FedDCA requires only about 30% encoded time overhead of FedDCA. Moreover, FedDCA achieves a substantial reduction in model size, which is less than 10% of FedDCA’s.
Privacy Analysis.
We evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks Carlini et al. (2021); Zhang et al. (2024a), which utilizes the autoregression nature of LLMXu et al. (2024). We analyze the domain inference attack in Section C.3.
We focus on one client’s instruction tuning in FedDIT, using FedDCA for instruction augmentation with 1k to 5k public instructions. We set up 10 clients with full participation for 10 rounds. Specifically, we record the average ROUGE-L score Lin (2004) denoted as $\rho$ for client $c_{0}$ in each round. The prompt to extract instructions is in Section C.1. For each setting, we repeat memory extraction 100 times and report $\rho$. Specifically, denote the generated $\mathcal{N}$ instructions as $\mathcal{I}$ and the client’s local instructions $\mathcal{I}^{l}$. The calculation is defined as $\rho=\frac{1}{\mathcal{N}}\sum_{i=1}^{\mathcal{N}}\texttt{ROUGE-L}(\mathcal{I}_{i},\mathcal{I}^{l})$.
(a) Different amounts of augmented public data for FedDIT.
(b) $\rho$ per round. Dotted lines represent the base-data-only setting.
Figure 4: Privacy analysis of memory extraction attacks. Data augmentation can effectively mitigate the privacy leakage risk.
Figure 4(a) shows no significant correlation between the public data ratio and privacy-preserving capability in the same training round. In addition, only using local data has a higher risk of privacy leakage than augmented methods. Additionally, Figure 4(b) shows $\rho$ trends per round. Initially, the $\rho$ for augmented settings increases, then decreases or converges, while the base-data-only scores continue to rise, especially in the code domain. This indicates that with more training rounds, base-data-only fine-tuning captures more privacy information, while the privacy leakage risk in augmented fine-tuning decreases or converges.
Scalability.
We scale the client number to 100 and 2 clients are randomly selected for FedDIT in each round. As the number of clients increases, the amount of local data on each client gradually grows. Table 4 shows that the model trained on only base data even outperforms Random Sampling in domains other than code and narrows the gap with FedDCA. However, because FedDCA aims to maximize the cross-client domain coverage, it achieves higher domain coverage and better performance.
Data Selection Scenarios.
We conduct the experiments on two settings: A) For each domain’s FedDIT, we randomly dispatch the in-domain data equally to each client. For example, in the code domain, as the size of the public dataset is 20,022, each client is assigned about 2,000 samples. Then we perform FedDIT on each client’s whole data. B) Following A’s setting, we first perform FedDCA to determine the client center set, and then we conduct the data selection based on the client center set through dense retrieval to select 200 samples for each client. Finally, we perform FedDIT on the selected data. Figure 6(b) demonstrates that FedDCA can achieve comparable performance to training with the full dataset by using only 10% data through data selection.
Held-out Setting.
If distributed clients aim to solve tasks based on existing knowledge, the public dataset will inevitably contain knowledge relevant to those domains. This could come from the original corpus (which can be converted into instruction-response pairs by GPT) or from pre-constructed instruction datasets on the website. So the distribution of the public dataset can be categorized as follows: containing held-in or held-out instructions. The held-in indicates that the public dataset contains instructions for the specific task that clients aim to solve, while the held-out indicates that the public dataset does not contain this task’s instructions. The paper’s default setting is the held-in setting.
Considering the held-out setting in the financial domain, given that the training set FinGPT and the test sets FPB, FiQA, and TFNS are all related to sentiment analysis tasks. We keep the setting of test sets and replace the FinGPT’s instructions in public data with data from a financial QA dataset (Sujet). The clients’ local data are still randomly sampled from FinGPT. This approach yields held-out public data.
Figures 6(d), 6(e) and 6(f) shows, FedDCA still achieves performance improvements compared to other baselines for the held-out setting. Additionally, using the Random Sampling data augmentation strategy resulted in performance degradation on the FiQA dataset. This further underscores the necessity of selecting an appropriate data augmentation strategy.
6.3 Ablation Study
Effect of Retrieval Number.
We report the performance and the corresponding domain coverage of FedDCA with different retrieval amounts on the four domains in Figure 5(b). We can see that the domain coverage of FedDCA is increasing along with the retrieval amount in different trends, as well as the performance except for the code domain. Noted that although the code domain is more about following a certain paradigm, which could perform well with a few data and more fine-tuning rounds, it still could benefit from the instruction augmentation.
Similarity Threshold.
Figure 5(a) shows the performance and the domain coverage of FedDCA w and w/o similarity threshold $\alpha$, where FedDCA w/o $\alpha$ is marked as FedDCA. Results show that FedDCA is slightly better than FedDCA in code, financial, and mathematical domains.
Impact of Different Cluster Number.
The hyperparameter $\xi$ is the number of clusters in the k-means algorithm. The experiment is conducted on $\xi=[N,2N,4N,8N]$, where $N$ is the number of clients. We report the domain coverage of the augmented dataset via FedDCA with different $\xi$ on the four domains in Figure 6(c). Results show that there is no best $\xi$ for all domains and $\xi=N$ is usually an acceptable choice.
+ FedDCA H-Eval M-Med FPB GSM8K
Random Sampling 32.51 (-0.42) 71.80 (+0.50) 63.92 (-0.27) 47.82 (+0.44)
Direct Retrieval 33.89 (-0.25) 72.10 (-0.10) 67.78 (+1.47) 52.53 (+1.66)
LESS 28.32 (+0.28) 70.80 (-0.20) 61.03 (+0.47) 43.92 (+0.79)
Self-Instruct 33.15 (+0.23) 70.70 (-1.20) 60.21 (+0.48) 50.25 (-0.54)
Table 3: Plug-and-Play. Performance (%) of FedDCA with different data augmentation methods. FedDCA can match full dataset performance using just 10% of the data through selection.
Metric Zero-shot Base Data Random Sampling FedDCA
H-Eval 29.88/- 34.14/0.8282 34.75/0.8685 35.92/0.9242
M-Med 70.60/- 72.40/0.8377 69.90/0.8497 73.30/0.9090
FPB 55.94/- 66.74/0.9339 61.05/0.9408 67.16/0.9800
GSM8K 23.27/- 49.12/0.8709 47.23/0.8812 50.26/0.9118
Table 4: Scalability. Performance (%) and domain coverage of FedDCA and other baselines. 100 clients with 2 clients per round. Results show the strong scalability of FedDCA.
(a) W and w/o $\alpha$
(b) Different Retrieval Amounts
Figure 5: Solid bars and lines are performance. Performance (%) and domain coverage of FedDCA (a) w and w/o similarity threshold $\alpha$; (b) w different retrieval amounts. Experiments show the necessity of similarity threshold for amplifying diversity and the strong applicability of FedDCA with various augmentation scale.
(a) Computational Cost
(b) Data Selection Scenarios
(c) Different $\xi$
(d) Held-out. FPB
(e) Held-out. FiQA
(f) Held-out. TFNS
Figure 6: (a) Computational cost comparison. (b) Domain coverage for different $\xi$. (c) Data selection scenarios. (d-f) Performance in the held-out setting. Results show the efficiency of FedDCA, the effectiveness of FedDCA in data selection scenarios, and the robustness of FedDCA in the held-out setting.
7 Conclusion
We reveal that in FedDIT, data heterogeneity has a non-monotonic relationship with model performance, while cross-client domain coverage significantly affects model effectiveness. To address this, we propose FedDCA, which optimizes domain coverage through greedy client center selection and retrieval-based instruction augmentation. The selection process ensures comprehensive cross-client domain coverage by greedily picking centers that represent maximal distinct regions of the instruction space. FedDCA further reduces client-side computation and enhances scalability using heterogeneous encoders. Extensive experiments comparing FedDCA with other nine baselines and across four domains demonstrate its effectiveness and plug-and-play capability.
Appendix
Appendix A Methodology
A.1 Complementary Description
Notation Table.
For convenience, we provide a detailed notation table in Table 5.
NOTATIONS REMARK
$N,C$ Clients number, client set $C=\{c_{1},c_{2},\ldots,c_{N}\}$.
$D^{p},D^{d},D^{c},D^{l}$ The public datasets, the in-domain data, the cross-client dataset, client’s local private data.
$N_{k}^{l},N_{k}^{p}$ Number of local private data on the $k$-th client, number of the retrieved public data on the $k$-th client.
$D_{k}^{l},D_{k}^{g},D_{k}$ The local private data on the $k$-th client, the retrieved public data on the $k$-th client, the augmented dataset on the $k$-th client.
$\Lambda,\mathcal{P},\mathcal{C}$ A specific sampling strategy that performs instruction augmentation on the server side, selected client center set, the cluster centers obtained locally and sent to the server for the greedy client center selection.
$\phi,\Delta\phi$ LLM’s pre-trained parameters, additional LoRA parameters.
$w,w_{enc},w_{p}$ Merged model parameters from the frozen LLM’s parameters $\phi$ and the additional LoRA parameters $\Delta\phi$, encoder model, projector model.
$F_{k}(w;\mathcal{D}),l(w;x,y)$ Loss of model $w$ over a specific dataset $\mathcal{D}$, the instructon tuning loss of model $w$ over a data sample $\left(x,y\right)$.
Table 5: Frequently used notation.
A.2 Direct Domain Data Retrieval
In this section, we first show the detail of instruction-based dense retrieval, which is both used in direct retrieval and FedDCA. Then, we explain the algorithm of the method direct retrieval.
A.2.1 Instruction Based Dense Retrieval
[…]
A.2.2 Direct Retrieval
The direct domain data retrieval only utilizes instructions of the client’s local data without responses to perform the retrieval-based domain data augmentation. The detailed algorithm is described in Algorithm 2. For each client, we start by encoding the local instructions using the encoder model $w_{enc}$. Next, we apply the k-means algorithm to cluster the embeddings into $\xi$ clusters. Then $\xi$ cluster centers are sent to the server for retrieval-based domain data augmentation. Subsequently, the retrieved public data is sent to the client for instruction tuning.
Algorithm 2 Direct domain data retrieval
1:Parameters:
2:Clients’ local datasets $D=\{D_{1},D_{2},\ldots,D_{N}\}$; Public dataset $D^{p}$; Local datasets $D^{l}=\{D^{l}_{1},D^{l}_{2},\ldots,D^{l}_{N}\}$; Number of clusters $\xi$; Encoder model $w_{enc}$; Pre-encoded public instruction embeddings $\mathcal{E}$.
3:$D^{g}\leftarrow\emptyset$
4:for $i\in\{1,2,\ldots,N\}$ do
5:     $\mathcal{E^{\prime}}\leftarrow\{w_{enc}(x)\mid x\in D_{i,instruction}^{l}\}$triangleright Encode the local instructions
6:     $\mathcal{C}\leftarrow\text{k-means}(\xi).fit(\mathcal{E^{\prime}})$triangleright Cluster local instructions, return cluster centers $\mathcal{C}$
7:     $\mathcal{S}\leftarrow\mathcal{C}\cdot\mathcal{E}^{T}$triangleright Compute the similarity score between $\mathcal{C}$ and $\mathcal{E}$
8:     $\mathcal{I}\leftarrow\emptyset$
9:     for $j\in\{1,2,\ldots,\xi\}$ do
10:         $\mathcal{S}^{\prime}\leftarrow\{s\mid s\in S_{j},\text{index}(s)\notin\mathcal{I}\}$triangleright Filter the selected indices
11:         $\mathcal{I^{\prime}}\leftarrow\text{indices of the top-}\frac{N_{k}^{p}}{\xi}\text{ elements in }\mathcal{S}^{\prime}$triangleright Retrieve the top $\frac{N_{k}^{p}}{\xi}$ indices
12:         $\mathcal{I}\leftarrow\mathcal{I}\cup\mathcal{I^{\prime}}$triangleright Update the selected indices
13:     end for
14:     $D^{g}_{i}\leftarrow\{D^{p}_{j}\mid j\in\mathcal{I}\}$triangleright Selected public instructions
15:     $D_{i}\leftarrow D^{g}_{i}\cup D_{i}^{l}$triangleright Obtain the augmented instruction dataset $D_{i}$
16:end for
Appendix B Related Works
Federated Instruction Tuning.
Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM Zhang et al. [2023c]; Wei et al. [2022]. Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios. So far, several FedIT frameworks Ye et al. [2024b, a]; Zhang et al. [2023b] have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods. While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigate the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
PrivateLoRA Wang et al. [2023b] addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data. While FewFedPIT Zhang et al. [2024b] focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
Non-IID data distribution is a common challenge in FL Karimireddy et al. [2020]; Li et al. [2020]; Sattler et al. [2019]; Li et al. [2019, 2021]. However, the impact of data heterogeneity on FedIT has not been fully explored. Shepherd Zhang et al. [2023b] constructing heterogeneity based on the topic of general instruction datasets. Results demonstrate that, unlike the consensus of traditional FL, for general FedIT, heterogeneity has a positive effect. By aggregating diverse instructions from clients, the diversity increases, thereby enhancing the model’s adaptability to various tasks. However, it just scratches the surface and does not explore issues in FedDIT. To fully understand the impact of data heterogeneity on FedDIT, we conduct a comprehensive analysis in this work and unveil a crucial finding: there is no monotonic correlation between the degree of non-independent and identically distributed (non-iid) and LLM’s performance in the context of FedDIT (Section 3).
Domain Instruction Augmentation.
In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code Nijkamp et al. [2023]; Luo et al. [2024], medical Zhang et al. [2023d], financial Yang et al. [2023b, a]; Zhang et al. [2023a]; Wu et al. [2023], mathematical Yue et al. [2024]; Luo et al. [2023]). Existing works tend to use open-source domain-specific instruction tuning datasets for training. However, the target domain may not always have corresponding ready-made domain-specific instruction datasets. Even if they exist, these datasets are often limited in scale.
Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets Wang et al. [2023a]; Zhang et al. [2023e]; Jiao et al. [2023]; Lee et al. [2024]; Xia et al. [2024]. For instance, Parrot Jiao et al. [2023] enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets. Furthermore, works like INSTA Lee et al. [2024] and LESS Xia et al. [2024] attempt efficient domain-specific instruction augmentation via dense retrieval. INSTA Lee et al. [2024] uses instructions without responses for effective retrieval. LESS Xia et al. [2024] assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval. 2) Using seed tasks for self-instruct Luo et al. [2024]; Wan et al. [2023]: Explore-Instruct Wan et al. [2023], for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct. 3) Scaling instructions from the web: Recent works Yue et al. [2024]; Zhou et al. [2024] highlight the immense potential of mining naturally occurring instructions from the internet. Compared to generated data, web-mined instructions exhibit less bias and greater diversity. MAmmoTH2 Yue et al. [2024] firstly retrieves domain-relevant texts and employs a LLM to extract Q-A pairs, further refining them into instruction-response pairs. Jiuzhang3.0 Zhou et al. [2024] distills GPT’s instruction generation capabilities into a smaller model and then uses it to generate instructions from the internet. In conclusion, models fine-tuned with augmented instructions have shown promising domain-specific capabilities.
To obtain a well-performing LLM in the specific domain within the distributed environment, we utilize a multi-domain dataset as public data on the server side and perform domain-specific instruction augmentation based on the client’s local instructions.
Appendix C Experiments
C.1 Implementation Details
We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round. Then, each available client performs FedDIT for 10 steps with AdamW optimizer, and the batch size is $B=32$ in a round. The initial learning rate is $55e-$ 5 with a cosine learning rate scheduler. Our experiment utilizes the widely used LLM, Llama3-8B [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B as the base model with 2048 max sequence length and adopts LoRA tuning method. The rank of LoRA is 16, and the scalar alpha is 16. For k-means Wu [2012], we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7. For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning Khosla et al. [2020]. We utilize bge-large-en-v1.5333https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions. While we utilize all-MiniLM-L6-v2444https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 as the client’s small encoder which outputs embeddings of 384 dimensions and bge-large-en-v1.5 as the server’s encoder for FedDCA.
Projector.
To reduce the computational overhead of clients, we propose a heterogeneous encoder method called FedDCA. We utilize a projector to perform the dimension alignment between the small encoder on the client side and the large decoder on the server side. As shown in Figure 7, where $d_{l}$ is the dimension of the client-side encoder, $d_{g}$ is the dimension of the server-side encoder.
Figure 7: The projector used on the server side for feature alignment, which consists of two fully connected layers and a ReLU activation layer in between.
Prompts Used in the Self-Instruct Data Generation.
To generate the Self-Instruct data, we prompt GPT-3.5 to generate the instruction with the designed prompt in Figure 8. Specifically, we randomly sample two examples from the client’s local data to guide GPT-3.5 generating the in-domain instruction and one example from the client’s local data for one-shot in-context learning to guide GPT-3.5 generating responses into the example’s format.
Figure 8: Prompts used in the Self-Instruct data generation. (a) Prompt for generating new instructions. Two examples are randomly sampled from the client’s local data for in-context demonstration. (b) Prompt for generating responses. We prompt GPT-3.5 to generate responses with a randomly selected example for one-shot in-context learning.
Prompt Used for Memory Extraction Attack.
As we use Llama3-8B as our base model and format the instructions and responses into the Alpaca’s format, to utilize the auto-regression nature of LLM to extract the instruction, we prompt the model to generate the instruction using the prompt in Figure 9, which is exactly the prefix of the Alpaca’s template.
Figure 9: Prompt used for memory extraction attack.
C.2 Baselines
Method Privacy Preserving API Cost Additional Information Performance Degradation Domain Coverage Oriented
FedAvg McMahan et al. [2017]
Random Sampling
LESS Xia et al. [2024]
Self-Instruct Wang et al. [2022]
Direct Retrieval
FedDCA (ours)
Table 6: Key differences between FedDCA and other baselines. FedDCA demonstrates the ability of: 1) privacy-preserving, 2) no API cost, 3) no additional information required, 4) avoiding performance degradation, and 5) aiming at domain coverage optimization. LESS requires additional information, as it needs access to the validation set for gradient-based retrieval.
In Table 6, we highlight five critical aspects of our approach compared with other methods Xia et al. [2024]; Wang et al. [2022]; Zhang et al. [2024b]. Our comparison focuses on five key dimensions:
  • Privacy Preserving: Most methods except Self-Instruct maintain data privacy, we analyze the privacy-preserving ability of FedDCA in Sections 5 and C.3.
  • API Cost: FedDCA requires no api calls. As Self-Instruct utilizes the GPT to generate instructions, it requires additional API calls, which incurs costs.
  • Additional Information: LESS needs validation set access for gradient-based retrieval, whereas FedDCA works with only the basic client’s local training data.
  • Performance Degradation: From Table 2 we see that there occurs performance degradation for FedAvg, Random Sampling, and Self-Instruct, while FedDCA avoids this issue.
  • Domain Coverage: Uniquely, FedDCA is specifically designed to optimize domain coverage for distributed environments, a feature absent in other baseline methods.
Specifically, Self-Instruct Wang et al. [2022] represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process. On the other hand, FewFedPIT Zhang et al. [2024b] attempts to augment local data by leveraging the pre-trained Llama2 model to perform self-instructed generation during training. However, two concerns exist: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation. 2) Generating instructions locally is very costly regarding both time and computational resources. Overall, generating instructions self-instructively is not a satisfactory method for the FedDIT scenario. LESS Xia et al. [2024] represents the augmentation methods through gradient feature retrieval. However, its performance is underwhelming. Compounding this issue, LESS presupposes access to the validation set, which is not always available, and necessitates an initial warmup on the public dataset before calculating gradients for both the public and validation data. This process becomes computationally burdensome with large public datasets.
In conclusion, these comparisons demonstrate FedDCA’s distinctive approach to balancing privacy preservation, computational overhead and domain coverage optimization, while maintaining minimal additional requirements.
C.3 Domain Inference Attacks
The server may inference the clients’ data domain when the domain data retrieval is performed on the server side. However, the proposed algorithm FedDCA is independent of the presence of a public dataset on the server. Even if the server does not have a public dataset, clients can upload their cluster centers to the server, which selects a set of client centers and sends them back to the clients. Each client can then retrieve data from the website based on the received client center by itself, thereby achieving data augmentation while maximizing the cross-client domain coverage.
In that case, since the server does not know the encoder used by the client, it cannot infer the semantic meaning of the embedding. Thus, for the server, it becomes significantly more challenging to infer the client’s domain, let alone apply any privacy protection techniques to the embeddings.
We provide two examples for illustration. Two different encoders are used as client’s and server’s respectively: BAAI/bge-large-en-v1.5 (denoted as $w_{1}$) and google-bert/bert-large-uncased (denoted as $w_{2}$). Both encoders output 1024-dimensional features.
Example 1: Both $w_{1}$ and $w_{2}$ take "hello world" as input, and the cosine similarity between their embeddings is 0.1829.
Example 2: Three instructions are used:
  • Instruction 1: Create an array of length 5 which contains all even numbers between 1 and 10.
  • Instruction 2: Write a replace method for a string class which replaces the given string with a given set of characters.
  • Instruction 3: What is the sentiment of this news? Please choose an answer from {negative/neutral/positive}. Teollisuuden Voima Oyj, the Finnish utility known as TVO, said it shortlisted Mitsubishi Heavy’s EU-APWR model along with reactors from Areva, Toshiba Corp., GE Hitachi Nuclear Energy, and Korea Hydro & Nuclear Power Co.
For these instructions, Instruction 1 is passed to $w_{1}$ and Instructions 2 and 3 are passed to $w_{2}$, which will result in three embeddings: $e_{1}$, $e_{2}$, and $e_{3}$. The cosine similarity between $e_{1}$ and $e_{2}$ is 0.1464, while the similarity between $e_{1}$ and $e_{3}$ is 0.1879. Instructions 1 and 2 are in the same domain, whereas they have a lower cosine similarity.
In conclusion, as demonstrated above, when the clients do not perform domain-specific instruction retrieval on the server side, the server cannot infer the client’s domain based on the uploaded embeddings.
C.4 How Far can FedDCA Go
Assume an extreme setting, where exists a pre-trained model that has already fine-tuned on the whole in-domain data of the internet or even all open-source public instruction datasets (it is actually unrealistic or temporary, as new data is constantly being generated), is there any room for instruction augmentation or FedDCA? To answer this question, we conduct the following experiments on four settings: A) Fine-tuning on the whole public dataset. B) Following A, perform further federated fine-tuning with only private clients’ local dataset. C) Following A, each client randomly samples 100 public data for fine-tuning. D) Following A, each client samples 100 public data through FedDCA and then performs FedDIT.
Setting H-Eval M-Med FPB FiQA TFNS GSM8K
A 38.41 69.10 75.33 41.09 71.60 53.75
B 43.90 64.30 77.97 66.54 78.26 56.40
C 41.46 66.90 74.50 29.81 70.51 57.01
D 44.12 70.90 81.43 72.00 78.81 58.90
Table 7: Performance (%) of setting A, B, C, and D in each domain.
As shown in Table 7, even with only 100 samples selected via FedDCA, the method not only prevents performance degradation but also helps the model achieve better generalization within the domain. This is attributed to exposure to more diverse in-domain data during training. Meanwhile, it can be observed that random sampling still leads to performance degradation compared to Setting A, and even performs worse than Setting B, where no data augmentation is applied. In the financial domain, performance degradation is also evident.
In conclusion, this experiment further highlights the necessity and effectiveness of FedDCA.
C.5 Augmentation Strategy Visualization
To more intuitively compare the domain coverage of different instruction augmentation methods, we randomly sample 5,000 instructions obtained through these methods and 10,000 in-domain instructions as the background, representing the distribution of specific domains in the public dataset. We then visualized the results using t-SNE van der Maaten and Hinton [2008], as shown in Figure 10. The plot shows that FedDCA encompasses most of the in-domain data, which is consistent with FedDCA’s domain coverage of each domain shown in Table 2. Also, we can observe that the random sampling strategy selects a lot of out-of-domain data while does not have good coverage in specific domains.
(a) Code
(b) Medical
(c) Financial
(d) Mathematical
Figure 10: Visualization of cross-client data distribution in different domains, performing t-SNE dimensionality reduction on retrieved instructions through various augmentation strategies. We randomly sample 10,000 in-domain samples as background while randomly sampling 5,000 samples from the cross-client augmented dataset for different instruction augmentation methods for comparison.